### v4

# Nova: Generative Language Models for Assembly Code with Hierarchical Attention and Contrastive Learning

Abstract
Binary code analysis is the foundation of crucial tasks in the security domain; thus building effective binary analysis techniques is more important than ever. Large language models (LLMs) although have brought impressive improvement to source code tasks, do not directly generalize to assembly code due to the unique challenges of assembly: (1) the low information density of assembly and (2) the diverse optimizations in assembly code. To overcome these challenges, this work proposes a hierarchical attention mechanism that builds attention summaries to capture the semantics more effectively, and designs contrastive learning objectives to train LLMs to learn assembly optimization. Equipped with these techniques, this work develops Nova, a generative LLM for assembly code. Nova outperforms existing techniques on binary code decompilation by up to 146.54%, and outperforms the latest binary code similarity detection techniques by up to 6.17%, showing promising abilities on both assembly generation and understanding tasks.
Nova: Generative Language Models for Assembly Code with
Hierarchical Attention and Contrastive Learning
Nan Jiang1, Chengxiao Wang1,2, Kevin Liu3, Xiangzhe Xu1, Lin Tan1, Xiangyu Zhang1 1Purdue University, 2University of Science and Technology of China, 3Lynbrook High School Correspondence: lintan@purdue.edu
1 Introduction
Binary code plays an irreplaceable role in the security domain, being the foundation of crucial tasks including vulnerability detection Güler et al. (2019); Duan et al. (2020); Chen et al. (2022b), malware detection Spensky et al. (2016); Aonzo et al. (2023); Xu et al. (2014), binary recovery Su et al. (2024); Zhang et al. (2021); Chen et al. (2022c), and legacy software maintenance Carbone et al. (2009); Carlini et al. (2015); Martin et al. (2010). For example, when performing tasks such as identifying attacks and malware, security analysts often only have access to assembly, i.e., the human-readable representation of binary code, which is extremely difficult to understand Su et al. (2024); Zhang et al. (2021); Chen et al. (2022c). Thus, combined with the increasing sophistication of cybercrime that poses significant threats worldwide (e.g., cybercrime is predicted to cost the world \$10.5 trillion annually by 2025 Sausalito (2020)), effective binary analysis techniques are in high demand.
Figure 1: Example that shows the semantics and diverse optimizations of assembly code.
Large language models pre-trained on source code have brought improvement in various software development domains Chen et al. (2022a); Liu et al. (2023); Chen et al. (2023); Le et al. (2022); Jiang et al. (2023); Xia et al. (2023). However, these LLMs are not designed for or trained with assembly corpus, not achieving their full potential on binary code analysis tasks such as binary code similarity Wang et al. (2022); Xu et al. (2023a), malware detection Su et al. (2024), and binary code decompilation Tan et al. (2024); Armengol-Estapé et al. (2024); Hosseini and Dolan-Gavitt (2022). Existing work applying LLMs on assembly code mainly piggybacks on encoder-style LLMs Wang et al. (2022); Su et al. (2024); Xu et al. (2023a), unable to benefit from the more extensive pre-training, updated architectures, scaling of state-of-the-art generative LLMs. Other work using generative LLMs for decompilation shows a low unit test passing rate of the decompiled programs Tan et al. (2024); Armengol-Estapé et al. (2024).
The challenges of leveraging generative LLMs for assembly code are twofold. First, compared to source code, assembly code has a lower information density. A short source-code sequence maps to a much longer assembly-code sequence that is often several times longer. Thus, assembly semantics span across a long sequence of tokens. Figure 1 (a) shows an example of a source code function that compares two integers, while Figure 1 (b) shows its corresponding assembly code optimized with O0 flag. In the O0-optimized assembly code, the five instructions from 10:   mov   -0x8(%rbp),%rax to 1c:   cmp   %eax,%edx perform the checking whether the value of x is smaller than the value of y (correspond to if (*(int*)x < *(int*)y) in the source code). A single assembly instruction alone represents little meaningful semantics in the source code. It is the combinations of many instructions and the dependencies between them represent the semantics. Such combinations of instructions are long, which is hard for LLMs to learn.
Second, assembly code is diverse due to compiler optimization. The assembly code of the same source code function looks dramatically different with different compiler optimization. Figure 1 (c) shows the assembly of the same function compiled with O1 and O0 flags, which consists of a significantly different set of instructions. Such syntax diversity is hard for LLMs to learn, preventing LLMs from obtaining consistently good performances on differently optimized assembly code.
In this work, we develop Nova, a generative foundation LLM pre-trained for assembly code with two key novelties. First, to address the low-information-density and long-sequence challenge, we design a hierarchical self-attention, which contains three categories of attention at different levels of granularity: intra-instruction attention, preceding-instruction attention, and inter-instruction attention. The key insight is to build attention summaries, i.e., we create per-statement attention labels, which act as the summary of a statement. We then use preceding-instruction attention to capture semantics between a token and its preceding instruction label and use inter-instruction attention for long dependencies. Besides, we design functionality contrastive learning and optimization contrastive learning objectives to train Nova to learn the semantics behind the diverse syntax of assembly.
This work makes the following contributions:
  • We propose a novel hierarchical attention mechanism that captures the assembly’s low-density semantics at three granularity levels.
  • We design contrastive learning objectives to train LLMs to learn assembly with diverse optimizations and encode assembly more efficiently.
  • We develop Nova, a generative foundation LLM with hierarchical attention and contrastive learning for assembly. Nova outperforms state-of-the-art (SOTA) on binary decompilation by up to 146.54% and on binary similarity detection by up to 6.17%.
  • We conduct a comprehensive analysis, illustrating the effectiveness of Nova’s novel designs: (1) Nova’s embeddings of assemblies successfully reflect code functionalities in the latent space, and (2) Nova’s hierarchical attention complements standard attention by learning different attention weight distributions, especially those reflecting long sequence semantics.
2 Approach
Figure 2 presents the overall approach of Nova. We build Nova on top of foundation models for source code Rozière et al. (2023); Li et al. (2023); Guo et al. (2024) to utilize their source code and natural language generation ability. We first collect large assembly corpora (Section 2.1). Section 2.2 describes Nova’s hierarchical attention. With the collected assembly corpora, we then pretrain Nova with language modeling and contrastive learning objectives (Section 2.3). Then, we fine-tune Nova on two important downstream tasks, binary code decompilation, and binary code similarity detection (Sections 2.4 and 2.5), to prove Nova’s effectiveness and benefits to the binary research domain.
Figure 2: Overview of developing Nova
2.1 Data Collection
We build our assembly data sets on top of existing source code datasets: The-Stack Li et al. (2023) and the AnghaBench da Silva et al. (2021). We compile the source code into executables with different optimization levels (i.e., O0, O1, O2 and O3), strip the executables to remove debug information, and disassemble them into assembly code. We treat every assembly function as a separate data sample. The breakdown statistics are in Table 1.
Datasets Source O0 O1 O2 O3 Total
AnghaBench 757.1K 743.1K 726.4K 718.7K 717.8K 3.7M
The-Stack 138.8K 125.1K 119.7K 116.9K 108.8K 609.3K
Table 1: Statistics (number of source code and assembly functions) of the pre-training datasets.
We perform certain normalization on the assembly functions: (1) removing all the “%” and comments, (2) adding whitespace around “,”, “(”, “)”, (3) converting all the hexadecimal numbers to decimal numbers, and (4) replacing the address of each instruction with special labels (e.g., replacing “0” and “4” in Figure 1 (b) with “[INST-1]” and “[INST-2]”) placing at the end of each instruction. More details are in Appendix A.1.
2.2 Hierarchical Self-Attention
Nova uses hierarchical self-attention that is specially designed to learn the low-information-density semantics in the long sequence of assembly code. Specifically, Nova learns the assembly code in an hierarchical way by providing a modified attention mask. Different from standard token-level attentions Vaswani et al. (2023); Radford and Narasimhan (2018); Radford et al. (2019); Brown et al. (2020), our hierarchical self-attention contains three categories at different levels of granularity.
Figure 3: Design of hierarchical attention
(1) Intra-Instruction Attention: Due to the low information density in assembly, intra-instruction attention is designed to capture the summary of every instruction, which is the standard causal attention but limited to tokens of each instruction (the yellow part in Figure 3). Tokens in different instructions have no attention weights. The “[INST]” label at the end of the instruction has attention to all the tokens in the instruction and thus captures the semantics of the entire instruction (e.g., “[INST-1]” captures the semantics of “mov eax, \$1”).
(2) Preceding-Instruction Attention In addition to the local semantics of each instruction, the use of assembly instructions (such as the choice of registers) depends on the context. For example, after the first instruction “mov eax, \$1”, the second instruction should not reuse “eax” to store another value “\$2” immediately. To capture such context, the preceding-instruction attention enables each token in an instruction to have attention to the “[INST]” label of the preceding instruction (the light green part in Figure 3).
(3) Inter-Instruction Attention To understand function semantics (i.e., functionality), which lies in the dependencies across different instructions, the inter-instruction attention is designed to let the “[INST]” label of each instruction have attention to all the labels of previous instructions. For example, “[INST-4]” has attention to “[INST-1]”, “[INST-2]”, and “[INST-3]” (the dark green part in Figure 3). The inter-instruction attention is only enabled for “[INST]” labels, as they represent the semantics of each instruction.
To sum up, the hierarchical self-attention breaks the semantics of assembly code into three parts. The intra-instruction attention captures the instruction summary, and the preceding-instruction attention captures the context with the preceding instruction. The inter-instruction attention learns the long dependencies across instructions on top of the “[INST]” labels that contain the instruction summary. Appendix A.2 shows how hierarchical self-attention works with text and source code.
Figure 4: Design of functionality and optimization contrastive learning (CL). “asm” denotes assembly.
2.3 Contrastive Learning
The syntax gap between assembly code and source code, and syntax diversity between differently-optimized assembly code make LLMs struggle to distinguish the semantics behind the syntax. Nova adopts contrastive learning technique Gao et al. (2021) during pre-training to train LLMs to encode assembly code meaningfully w.r.t semantics.
The standard pre-training objective is language modeling by minimizing the negative likelihood of code in the pre-training corpus Radford and Narasimhan (2018), notated as $L_{lm}$. In addition, Nova is pre-trained with two new objectives, $L_{fcl}$ for functionality contrastive learning and $L_{ocl}$ for optimization contrastive learning.
Functionality CL Functionality CL trains Nova to focus more on the functionalities of assembly code rather than the syntax. Code with the same functionality (assemblies from the same source code), should be encoded closer in the latent space. For instance, in Figure 4 (a), embeddings of source and assembly code of function “cmp” are closer to each other, and the same for function “sort”.
Let $e^{s}_{f}$ be the embedding of function $f$ in $s$ form ($s=-1$ for source code, and $s\in[0,1,2,3]$ for O0 to O3 optimized assembly). For simplicity, let $S=[-1,0,1,2,3]$ be the domain of $s$. We use the average of all the “[INST]” tokens’ embedding as the embedding of the whole assembly function, as each “[INST]” token is supposed to capture the semantics of that instruction by the design of hierarchical self-attention. Functionality CL optimizes Nova with the constraint:
[…]
Such constraints can be trained by optimizing the embeddings of a batch of functions, each function in two different forms. For the example in Figure 4 (b), there are two forms (source code and O0 assembly) of four functions. Once Nova encodes the batch of source code and assembly functions, we calculate the distance matrix $\{D_{ij}\}_{f_{i},f_{j}\in F}=\{d(e^{s}_{f_{i}},e^{t}_{f_{j}})\}$, and minimize the loss:
$L_{fcl}=-\mathop{log}\sum\limits_{s,t\in S}\sum\limits_{f_{i}\in F}\left(1-\frac{\mathop{exp}\left(d(e^{s}_{f_{i}},e^{t}_{f_{i}})\right)}{\sum_{f_{j}\in F}\mathop{exp}\left(d(e^{s}_{f_{i}},e^{t}_{f_{j}})\right)}\right)$
This objective minimizes the distance between embeddings for the same function, i.e., the diagonal in the distance matrix.
Optimization CL LLMs can be confused if being asked to directly connect a source code function to its O3-optimized assembly, due to their dramatically different syntax. Such a huge gap can be filled by learning how the source code is transformed to O0, O1, O2 and eventually to O3 assembly, as the optimization levels are ordered.
Higher-level optimization applies a super-set of optimization rules compared to lower-level optimization. Nova learns such order with the optimization CL objective, encoding differently-optimized assembly code orderly. Optimization CL optimizes Nova with the constraint:
$\forall f\in F,\forall s<t_{1}<t_{2}\in S,\;d(e^{s}_{f},e^{t_{1}}_{f})\leq d(e^{s}_{f},e^{t_{2}}_{f})$
Intuitively, this ensures that the more optimizations applied, the larger the difference between embeddings of optimized and unoptimized code. For instance, Figure 4 (c) and (d) illustrate that for the same function “cmp”, the distance between source code and assembly increases when the optimization level increases. Formally, optimization CL minimizes the following loss:
$L_{ocl}=\sum\limits_{f\in F}\sum\limits_{s<t_{1}<t_{2}\in S}\mathop{max}\left(0,d(e^{s}_{f},e^{t_{1}}_{f})-d(e^{s}_{f},e^{t_{2}}_{f})\right)$
Overall, the final training loss combines the three: $L=L_{lm}+\lambda(L_{fcl}+L_{ocl})$, where $\lambda$ is set to 0.1 to balance the losses in this work.
2.4 Task 1: Binary Code Decompilation
[…]
2.5 Task 2: Binary Code Similarity Detection
Binary code similarity detection (BCSD) aims to measure the similarity between two binary code snippets Wang et al. (2022); Su et al. (2024), which is the foundation of various applications such as plagiarism detection Luo et al. (2014); Sæbjørnsen et al. (2009) and vulnerability detection David and Yahav (2014); David et al. (2018, 2017, 2016).
A widely used setting is taking a query assembly of the function $f^{q}$ that is compiled with one optimization level (denoted by $s$), and a pool of candidate assembly of $K$ functions (notated by $f^{p}_{i},$ $1\leq i\leq K$) compiled with a different optimization level (denoted by $t\neq s$). There exists a unique candidate assembly coming from the same source code as the query ($\exists!1\leq i\leq K,$ $f^{p}_{i}=f^{q}$, called the positive candidate). Nova is fine-tuned to encode these binaries, so that the positive candidate has the highest similarity with the query assembly among the pool. The learning objective is:
[…]
, where we follow previous work Su et al. (2024) to let $s$ be O0-assembly and $t$ be O3-assembly, which is the hardest setting.
3 Experimental Setup
3.1 Pre-Training
We use the data collected from AnghaBench and The-Stack for pre-training. We pre-train Nova starting from DeepSeek-Coder Guo et al. (2024), and the hierarchical attention is applied on half of the attention heads to balance between its effectiveness and the existing knowledge in the standard attention layers. Nova is pre-trained with language modeling for one epoch, followed by contrastive learning objectives for another epoch.
3.2 Binary Code Decompilation
Training Data We sample 2.16M assembly-to-source-code pairs (0.338B tokens) from the pre-training corpus to build the BCD fine-tuning data.
Test Data We use HumanEval-Decompile Tan et al. (2024) as the test benchmark, which was not used in training. HumanEval-Decompile is derived from the C language adaptation of the HumanEval Chen et al. (2021) benchmark and provides test cases in evaluating functionality correctness. HumanEval-Decompile contains 164 C functions, each compiled with O0O3 optimization flags and disassembled into X86-64 assembly.
Baselines Nova is compared with GPT-3.5, GPT-4, and the previous SOTA LLM4Decompile Tan et al. (2024). LLM4Decompile trains DeepSeek-Coder using the same AnghaBench corpus, and it is the first LLM-based technique that aims to generate executable decompilations.
Evaluation GPT-3.5 and GPT-4 are prompted with three-shot examples, while LLM4Decompile and Nova samples 20 decompilations per assembly, using the temperature of 0.2 and top_p of 0.95 Chen et al. (2021). The generated decompilations are executed against the test cases and both Pass@1 and Pass@10 Chen et al. (2021) are reported.
3.3 Binary Code Similarity Detection
Training Data To compare Nova with existing works on BCSD fairly Wang et al. (2022); Su et al. (2024), we use BinaryCorp-3M Wang et al. (2022) as the fine-tuning data for BCSD, which contains the O0 and O3 assembly of 224,606 functions.
Test Data Following existing work Su et al. (2024); Xu et al. (2023a), we use real-world benchmarks, Binutils, Curl, ImageMagick, SQLite, OpenSSL, and Putty, as the test benchmarks, which are nonexistent in the training data.
Baselines Nova is compared with jTrans Wang et al. (2022), DiEmph Xu et al. (2023a) and CodeArt Su et al. (2024). jTrans is a Transformer Vaswani et al. (2023) encoder trained on binaries with masked token prediction and jump target prediction tasks. DiEmph uses an instruction deemphasis technique to prevent the model from learning instruction distribution biases introduced by compilers. CodeArt proposes a regularized attention mask for encoder models to capture instructional semantics and data dependencies.
Evaluation We randomly sample $K$ source code functions from each project, compile them into binaries with O0 and O3 optimization flags, and disassemble them into X86-64 assemblies. BCSD techniques encode these assemblies into embeddings (Nova uses the average embeddings of all the “[INST]” tokens in an assembly as its embedding). Then each O0 assembly is used as the query to calculate their similarity with the $K$ O3 candidate assemblies (using cosine similarity). Metric Recall@1 is the ratio of queries for which the candidate from the same source code has the highest similarity among all the candidates.
Appendix A.3 contains additional details such as training hyper-parameters.
4 Results
4.1 Binary Code Decompilation
4.1.1 Comparison with SOTA Techniques
Table 2 shows the Pass@1 of the decompiled code from assemblies on the HumanEval-Decompile benchmark. The results are grouped by optimization level (i.e., the benchmark contains 164 assemblies of each optimization level to decompile), and the average is also reported.
Optimization O0 O1 O2 O3 Avg.
GPT-3.5 6.80 5.64 4.36 3.93 5.18
GPT-4 17.77 15.12 12.65 11.25 14.20
LLM4Decompile-1B 12.26 7.22 8.38 7.96 8.96
Nova-1B 31.19 17.29 18.72 15.58 22.09
LLM4Decompile-6B 23.01 15.95 16.95 14.79 17.68
Nova-6B 42.07 28.04 25.00 22.56 29.42
Table 2: Pass@1 on HumanEval-Decompile.
Optimization O0 O1 O2 O3 Avg.
GPT-3.5 8.95 7.77 5.93 5.12 6.94
GPT-4 25.64 20.65 18.70 18.03 20.76
LLM4Decompile-1B 17.95 12.05 13.90 12.51 14.10
Nova-1B 41.11 29.81 31.18 26.24 32.09
LLM4Decompile-6B 33.77 24.25 23.94 23.81 26.44
Nova-6B 51.06 37.83 35.79 34.63 39.83
Table 3: Pass@10 on HumanEval-Decompile.
Overall, Nova’s Pass@1 is higher than all SOTA binary decompilation techniques and general LLMs GPT-4 and GPT-3.5, which are orders of magnitude larger than Nova. Specifically, for each optimization level, Nova consistently decompiles more assemblies into source code correctly than LLM4Decompile, GPT-3.5, and GPT-4. With the same model size, Nova-1B outperforms LLM4Decompile-1B by 146.54%, i.e., averaged Pass@1 of 22.09% versus 8.96%. Nova-6B outperforms LLM4Decompile-6B by 66.40%: the averaged Pass@1 is 29.42% versus 17.68%.
Table 3 shows that Nova still outperforms SOTA techniques with a significant margin under the measurement of Pass@10. Examples of Nova’s correct decompilation are provided in Appendix A.4.
4.1.2 Ablation Study
We conduct an ablation study by comparing Nova-1B with the following models:
  • Nova-CL-HA: Removing contrastive learning and hierarchical self-attention. This is simply training DeepSeek-Coder-1.3B on the assembly corpus using language modeling.
  • Nova-HA: Removing the hierarchical self-attention, training DeepSeek-Coder-1.3B on the assembly corpus using both the language modeling and contrastive learning objectives.
Nova-CL-HA can be viewed as our reproduction (retrain) of LLM4Decompile-1B.
Optimization O0 O1 O2 O3 Avg.
LLM4Decompile-1B 17.95 12.05 13.90 12.51 14.10
Nova-CL-HA 17.80 13.32 13.26 10.03 13.60
Nova-HA 25.12 15.64 16.07 12.71 17.39
Nova 31.19 17.29 18.72 15.58 22.09
Table 4: Ablation study of Nova-1B (Pass@1).
Table 4 shows the results of the ablation study, reported by the Pass@1 on HumanEval-Decompile. Nova-CL-HA shows comparable Pass@1, which we considered as variance in reproducing the same approach. With additional contrastive learning objectives, Nova-HA improves the Pass@1 on all optimization levels over Nova-CL-HA, showing a 27.87% higher averaged Pass@1. Further applying the hierarchical self-attention boosts the overall Pass@1 from 17.39% to 22.09%.
Benchmarks jTrans DiEmph CodeArt Nova-1B Nova-6B
Binutils 0.68 0.80 0.84 0.87 0.89
Curl 0.72 0.84 0.86 0.89 0.94
ImageMagick 0.53 0.71 0.78 0.86 0.90
SQLite 0.73 0.79 0.78 0.77 0.78
OpenSSL 0.70 0.83 0.88 0.90 0.92
Putty 0.63 0.72 0.69 0.72 0.71
Avg. 0.67 0.78 0.81 0.84 0.86
Table 5: Recall@1 on benchmarks with $K=50$.
Benchmarks jTrans DiEmph CodeArt Nova-1B Nova-6B
Binutils 0.60 0.63 0.81 0.79 0.79
Curl 0.63 0.80 0.82 0.86 0.88
ImageMagick 0.54 0.71 0.76 0.79 0.81
SQLite 0.62 0.72 0.74 0.73 0.72
OpenSSL 0.60 0.80 0.87 0.88 0.90
Putty 0.58 0.64 0.64 0.65 0.64
Avg. 0.60 0.72 0.77 0.78 0.79
Table 6: Recall@1 on benchmarks with $K=100$.
Benchmarks jTrans DiEmph CodeArt Nova-1B Nova-6B
Binutils 0.51 0.64 0.74 0.73 0.73
Curl 0.57 0.77 0.78 0.83 0.84
ImageMagick 0.39 0.51 0.67 0.73 0.75
SQLite 0.56 0.65 0.68 0.68 0.69
OpenSSL 0.54 0.71 0.82 0.84 0.88
Putty 0.49 0.58 0.55 0.55 0.58
Avg. 0.51 0.64 0.71 0.73 0.75
Table 7: Recall@1 on benchmarks with $K=200$.
Benchmarks jTrans DiEmph CodeArt Nova-1B Nova-6B
Binutils 0.40 0.57 0.70 0.65 0.67
Curl 0.43 0.62 0.69 0.73 0.76
ImageMagick 0.25 0.42 0.58 0.61 0.65
SQLite 0.43 0.59 0.62 0.59 0.62
OpenSSL 0.43 0.61 0.76 0.78 0.82
Putty 0.38 0.50 0.49 0.47 0.51
Avg. 0.39 0.55 0.64 0.64 0.67
Table 8: Recall@1 on benchmarks with $K=500$.
4.2 Binary Code Similarity Detection
4.2.1 Comparison with SOTA Techniques
Tables 8,  8,  8 and  8 show the Recall@1 of Nova and existing BCSD techniques with pool size $K$ of 50, 100, 200 and 500 on the six benchmarks. Underlined numbers indicates the best in each benchmark, while wavey underlined numbers denote the tied best (we only mark Nova-1B for clearer illustration).
Overall, Tables 888 and 8 show that on average, Nova-1B and Nova-6B achieve the highest Recall@1 (in bold) under all four settings of $K$. Nova-6B further outperforms Nova-1B and achieves the highest averaged Recall@1 under all four settings, ranking the ground-truth of 5%, 2%, 4%, and 3% more queries the most similar correspondingly compared to CodeArt.
Nova-1B consistently outperforms existing techniques with higher Recall@1 when $K$ is 50, 100, and 200, meaning it correctly ranks ground-truth of 3%, 1%, and 2% more queries as the most similar. Under the setting of $K=500$, Nova-1B ties with CodeArt with the same highest Recall@1. When looking into each individual benchmark, Nova-1B always wins on the most benchmarks under different settings of pool size $K$. For instance, Nova-1B wins on four benchmarks while DiEmph only wins on SQLite when $K=50$.
$K$ Nova-CL-HA Nova-HA Nova
50 0.81 0.83 (+0.02) 0.84 (+0.01)
100 0.76 0.78 (+0.02) 0.78
200 0.70 0.70 0.73 (+0.03)
500 0.60 0.62 (+0.02) 0.64 (+0.02)
Table 9: Ablation study of Nova-1B (Recall@1)
4.2.2 Ablation Study
Table 9 shows the averaged Recall@1 of Nova-CL-HA, Nova-HA (same as in Section 4.1.2), and Nova-1B under four pool size settings. With contrastive learning objectives, Nova-HA improves Nova-CL-HA under three settings ($K=50,100,200$) with 2% higher Recall@1. With hierarchical attention, Nova further outperforms Nova-HA under three settings ($K=50,200,500$). Detailed ablation study results on each benchmark are provided in Appendix A.5.
4.3 Analytic Experiments
4.3.1 How are Nova’s embeddings better?
[…]
Figure 5 (b) shows that Nova encodes assemblies into clusters of functionalities. The assemblies for the same functionality (i.e., the same task) are encoded closer to each other than Nova-CL-HA does in Figure 5 (b). The results show that our hierarchical attention and contrastive learning techniques effectively group codes of similar functionalities together for better assembly foundation models. Embedding of Nova-HA and additional quantitative analysis are shown in Appendix A.6.
Figure 5: PCA of embeddings calculated by Nova-CL-HA and Nova, for HumanEval-Decompile assemblies.
4.3.2 What does hierarchical attention learn?
We conduct quantitative analysis on the attention weights produced by different models.
Figure 6: Quantitative analysis of attention weights.
Entropy Figure 6 (a) shows the entropy of attention-weight distributions in each layer. We separate the attention heads as standard attention (green line) and hierarchical attention (red line), since Nova applies hierarchical attention to half of the attention heads in each layer (Section 3.1).
Nova’s hierarchical attention heads produce significantly lower entropy, suggesting its attention layer is more confident in learning specific relationships than the other models’ attention layers. The standard attention heads in Nova show patterns similar to those of Nova-CL-HA and Nova-HA, allowing Nova learning the standard attention to capture the “soft” relationship between possible tokens pairs. The result suggests that hierarchical attention complements standard attention with additional knowledge.
[INST] Token Figure 6 (b) shows the attention weights paid to the “[INST]” tokens. Nova’s hierarchical attention heads pay more attention to the “[INST]” tokens than standard attention, which may be because these “[INST]” tokens contain instruction summary and long dependencies and thus are more informative. Additional analysis and examples are given in Appendix A.7.
5 Related Work
5.1 Binary Models
Machine learning models are widely used in binary program analysis tasks. However, these models are typically designed for specific downstream tasks such as binary code similarity detection Pei et al. (2020); Xu et al. (2023a); Wang et al. (2022), variable name prediction Chen et al. (2022c); Xu et al. (2023b); Zhang et al. (2021), binary code type inference Pei et al. (2021), and are built from scratch.
Recent techniques have started to pre-train foundation LLMs for binaries. CodeArt Su et al. (2024) pre-trains encoder-style LLMs with a regularized attention design to better encode assembly code semantics, showing good accuracy on binary code understanding tasks (e.g., binary code similarity detection and malware family classification). SLaDe Armengol-Estapé et al. (2024) trains BART Lewis et al. (2019) models on assembly, and LLM4Decompile Tan et al. (2024) trains DeepSeek-Coder with assembly for binary code decompilation. However, CodeArt does not generalize to generation tasks due to its encoder architecture. SLaDe and LLM4Decompile are limited in performance due to a lack of special designs for assembly. In contrast, Nova addresses both limitations, by using the proposed hierarchical attention and contrastive learning objectives, outperforming existing techniques on both understanding (binary code similarity detection) and generation (binary code decompilation) tasks.
5.2 Large Source-Code Models
LLMs demonstrate promising results on many code-related tasks, such as code generation Chen et al. (2022a); Liu et al. (2023); Chen et al. (2023); Le et al. (2022); Yue et al. (2021); Chen et al. (2021); Nijkamp et al. (2022); Fried et al. (2023); Rozière et al. (2023); Guo et al. (2024), bug fixing Jiang et al. (2023); Xia et al. (2023) and vulnerability fixing Wu et al. (2023); Steenhoek et al. (2023); He and Vechev (2023). The advances in using LLMs are attributed to the knowledge learned from massive source code and natural language text in their training datasets Touvron et al. (2023); OpenAI (2023). Nova is designed and trained for assembly, which has unique challenges such as low information density and diverse optimization.
6 Conclusion
This work develops Nova, a generative foundation LLM for assembly code, which incorporates two key novelties (hierarchical attention and contrastive learning objectives) to address the unique challenges of assembly code. Evaluation on downstream tasks shows the effectiveness of Nova, which outperforms existing techniques on binary code decompilation by up to 146.54% and outperforms the latest binary code similarity detection techniques by up to 6.17%. We expect our hierarchical attention and contrastive learning techniques to benefit source code and natural language foundation models, which remains as future work.
7 Limitations
One limitation is that Nova is X86-specific, as we only collect X86 assembly corpus for pre-training. This design choice is mainly affected by two facts: (1) X86 assembly is used and explored in a wide range of binary tasks Wang et al. (2022); Su et al. (2024); Xu et al. (2023a); Chen et al. (2022c) compared to other assembly languages, and (2) computation limitations. However, the proposed techniques are independent of X86 assembly. Low information density and compiler optimization are the common challenges of most assembly languages such as X86, ARM, and MIPS. The proposed techniques can be applied to the future development of multi-lingual assembly LLMs.
Another potential limitation is the scale of models. We develop Nova-1B and Nova-6B. These two sized LLMs show impressive ability in assembly code decompilation and encoding. There might be potential benefit of developing larger Nova models. However, due to the computing resources limitation, we are unable to explore that in this work.
ACKNOWLEDGMENTS
This research was supported in part by NSF 1901242 and 2006688 and a CFI fund.
Appendix A Appendix
A.1 Data Collection
This section provides additional details of the data collection. To collect assemblies from The-Stack, we attempt to compile 4 million C programs, of which 138.8K is compiled successfully. We do not collect more due to the computation resource limitations.
For the 757.1K and 138.8K source code that successfully compiled into executables (using gcc) from AnghaBench and The-Stack, we disassemble them using objdump. objdump was not able to successfully disassemble all the executables, resulting in some empty assembly code. Thus, the number of O0O1 we obtain from each corpus is different and smaller than the number of source codes as shown in Table 1.
Figure 7 shows an example of preprocessing the raw assembly code as described in Section 2.1.
Figure 7: Example of assembly code preprocessing
A.2 Hierarchical Self-Attention
The hierarchical self-attention is designed for assembly code, yet the input to LLMs may still contain text or source code. Figure 8 illustrates how the hierarchical attention works with text or source code in the input. As existing LLMs have shown good performance on text and source code using the standard self-attention, we keep the standard causal attention mask within and between any chunks of text or source code in the input (the light grey part shown in Figure 8).
The attention from text or source code to assembly code (and vice versa) is only allowed for the “[INST]” tokens as they are supposed to contain the assembly instruction summaries.
Figure 8: Hierarchical attention with text input.
Benchmarks Nova-CL-HA Nova-HA Nova-1B
Binutils 0.86 0.88 0.87
Curl 0.84 0.87 0.89
ImageMagick 0.79 0.80 0.86
SQLite 0.80 0.83 0.77
OpenSSL 0.90 0.92 0.90
Putty 0.68 0.66 0.72
Avg. 0.81 0.83 0.84
Table 10: Ablation study with $K=50$.
Benchmarks Nova-CL-HA Nova-HA Nova-1B
Binutils 0.80 0.82 0.79
Curl 0.84 0.84 0.86
ImageMagick 0.70 0.72 0.79
SQLite 0.74 0.78 0.73
OpenSSL 0.89 0.89 0.88
Putty 0.59 0.60 0.65
Avg. 0.76 0.78 0.78
Table 11: Ablation study with $K=100$.
Benchmarks Nova-CL-HA Nova-HA Nova-1B
Binutils 0.71 0.74 0.73
Curl 0.80 0.73 0.83
ImageMagick 0.61 0.63 0.73
SQLite 0.68 0.71 0.68
OpenSSL 0.85 0.87 0.84
Putty 0.53 0.53 0.55
Avg. 0.70 0.70 0.73
Table 12: Ablation study with $K=200$.
Benchmarks Nova-CL-HA Nova-HA Nova-1B
Binutils 0.62 0.65 0.65
Curl 0.67 0.71 0.73
ImageMagick 0.46 0.51 0.61
SQLite 0.61 0.62 0.59
OpenSSL 0.77 0.79 0.78
Putty 0.46 0.46 0.47
Avg. 0.60 0.62 0.64
Table 13: Ablation study with $K=500$.
A.3 Training Details
This section provides additional details of training. We pre-train Nova starting from DeepSeek-Coder, using the language modeling objective ($L_{lm}$) for one epoch on the AnghaBench and The-Stack corpora. This step uses a batch size of 128, with the input truncated by a 1,024 tokens limit. The model weights are updated using the AdamW optimizer. The learning rate is $5e^{-5}$, using 1000 steps of warm-up and a cosine decay to adjust the learning rate.
Then, the model is further pre-trained with the combination of language modeling and contrastive learning objectives ($L=L_{lm}+\lambda(L_{fcl}+L_{ocl})$), with $\lambda$ set to 0.1. To train with the functionality contrastive learning objective, we discard any source code that misses any one of O0O3 assemblies and also discard the source code whose O2 assembly is the same as its O3 assembly. As a result, this step is only trained for 0.36M data samples for one epoch. The batch size is 64, with the input truncated by a 1,024 tokens limit. The learning rate is $2e^{-5}$ using the AdamW optimizer.
The fine-tuning of both BCD and BCSD uses a batch size of 64, with the input truncated by a 2,048 token limit. Similarly, the learning rate is $2e^{-5}$ using the AdamW optimizer, and the model is fine-tuned for one epoch.
Infrastructure The training are conducted on eight NVIDIA RTX A100 GPUs, each with 40GB memory. Our implementation is based on Huggingface’s implementation of DeepSeek-Coder  [^1] https://huggingface.co/deepseek-ai/deepseek-coder-1.3b-base, PyTorch  [^2] https://pytorch.org/get-started/locally/, and DeepSpeed  [^3] https://github.com/microsoft/DeepSpeed.
Figure 9: Nova-1B correctly decompiles HumanEval-Decompile task 0, while GPT-4 and LLM4Decompile-1B fail.
Figure 10: Nova-1B correctly decompiles HumanEval-Decompile task 20, while LLM4Decompile-1B fail.
Figure 11: PCA of embeddings calculated by Nova-CL-HA, Nova-HA, and Nova.
Figure 12: t-SNE of embeddings cauclated by Nova-CL-HA, Nova-HA, and Nova.
A.4 Binary Code Decompilation Case Studies
Figure 9 shows an example from HumanEval-Decompile (task_id 0). Given the O1-optimized assembly code, GPT-4 fails to figure out the number of function arguments correctly, missing one important argument “float e”, and thus produces wrong functionality in the decompiled code. LLM4Decompile-1B makes similar mistakes and also misses the inner nested for loop. Nova-1B correctly decompiles the assembly into source code, where the ground truth is checking if any two elements in the given list *a (with size n) are close to each other than a given threshold e.
Figure 10 shows another more complex example, HumanEval-Decompile task_id 20. Nova-1B correctly decompiles the source code, successfully figuring that the function is trying to find the two elements that are closest to each other in the given array *a, with the result stored in minmax.
A.5 Binary Code Similarity Detection Ablation Study
Table $13$ show the detailed ablation study results of BCSD. Nova wins on the most benchmarks when $K=100$ or $500$, and ties with Nova-HA when $K=50$, or $200$.
A.6 Additional Analysis of Embedding
Additional Embedding Visualization
Figure 11 shows the full results of PCA of embeddings provided by Nova-CL-HA, Nova-HA, and Nova, on randomly sampled seven examples. Compared with Nova-CL-HA, Nova-HA including contrastive learning objectives in the pre-training, can separate the embeddings of assemblies with different functionalities better. Nova-HA clearly encode “Task 143” (orange points) away from the others. Nova’s embeddings group the assemblies by functionalities more precisely than Nova-HA, suggesting that hierarchical attention enhances the training of contrastive learning objectives to learn more effective encoding.
Figure 12 shows the results using another dimensionality reduction technique, t-SNE van der Maaten and Hinton (2008), where Nova’s embeddings are consistently more distinguishable by functionalities.
Figure 13: Distances between embeddings of different-formats-same-functionality code and same-format-different-functionalities code.
Figure 14: Comparison of attention distribution among standard and hierarchical heads in the final layer.
Quantitative Analysis
We also perform quantitative analysis to further support the conclusion from visualization. Figure 13 (a) shows the distribution of $l_{2}$ distances between different-formats-same-functionality code in HumanEval-Decompile (“format” is defined as one of source code, O0, O1, O2, or O3 assembly). The figure shows five distributions, (src, O0), (src, O1), (src, O2), and (src, O3). Each distribution calculates the distances between embeddings of two formats of code, e.g., src, O0 refers to the source code and O0 assembly for the same task in HumanEval-Decompile.
Figure 13 (a) shows that Nova’s embeddings for same-functionality-different-formats code are closer to each other in the latent space. For the same task, all the O0O3 assemblies’ embeddings have a smaller $l_{2}$ distance to the source code embedding compared to that of Nova-CL-HA’s and Nova-HA’s embeddings. Another interesting finding is that for Nova, the distances between assembly and source code increase with the optimization level of assemblies increases. This trend matches what the optimization CL tries to optimize for.
Figure 13 (b) shows the distribution of $l_{2}$ distance between same-format-different-functionalities code from the HumanEval-Decompile benchmark. With the same format (source code or the same optimized assembly), Nova encodes the assemblies with different functionalities farther away from each other compared to Nova-CL-HA. Nova-CL-HA’s embeddings cannot significantly reflect the functionality differences between assemblies, while Nova’s embeddings can.
Nova-HA’s embeddings show high $l_{2}$ distance in both Figure 13 (a) and (b), suggesting that Nova-HA tries to decentralize all the code in the embedding space even if they have the same functionality, which is not as desired as Nova.
A.7 Additional Analysis of Attention
Figure 15: Learned per-instruction soft attention observed in the lower layers
Figure 14 shows the visualizations of attention weights in the final transformer layer of two select heads with standard attention and two heads with learned hierarchical attention. Standard attention exhibits two typical patterns, namely diagonal attention (i.e. tokens attending to themselves or nearby tokens, shown in Figure 14 (a)), and broad attention (i.e. a single token attending broadly to the entire sequence, shown in Figure 14 (b)). In contrast, in Nova’s hierarchical attention, attention weights are allocated among distinct segments, each corresponding to an instruction (shown in Figure 14 (c)), that focus on tokens comprising that instruction (e.g. opcodes and operands, shown in Figure 14 (d), attentions are paid to “push”, “mov”, etc.).
Quantitatively, we have determined broad attention accounts for as much as 30% of all attention in standard heads, especially in layers 1-8 (consistent with the findings of (Clark et al., 2019)), whereas in Nova’s hierarchical attention, no more than 5% or all attention is allocated to each instruction segment. This validates our goal of learning instruction-aware hierarchical attention in Nova.
In addition, in lower layers, we have observed attention weights to be softly distributed among tokens comprising each instruction (Figure 15), which suggests Nova initially models cross-relations among operation codes and operands in the first few layers, and later pools their summary representation into the [INST] token in the later layers. This is also supported by Figure 6, where the Nova’s hierarchical attention (red line) shows a decreasing trend of entropy. This means the hierarchical attention is softer in lower layers, and becomes focused in higher layers.

### v5

@@
 Equipped with these techniques, this work develops Nova, a generative LLM for assembly code.
-Nova outperforms existing techniques on binary code decompilation by up to 146.54%, and outperforms the latest binary code similarity detection techniques by up to 6.17%, showing promising abilities on both assembly generation and understanding tasks.
-Nova: Generative Language Models for Assembly Code with
-Hierarchical Attention and Contrastive Learning
-Nan Jiang1, Chengxiao Wang1,2, Kevin Liu3, Xiangzhe Xu1, Lin Tan1, Xiangyu Zhang1 1Purdue University, 2University of Science and Technology of China, 3Lynbrook High School Correspondence: lintan@purdue.edu
+Nova outperforms existing techniques on binary code decompilation by up to 14.84 – 21.58% (absolute percentage point improvement) higher Pass@1 and Pass@10, and outperforms the latest binary code similarity detection techniques by up to 6.17% Recall@1, showing promising abilities on both assembly generation and understanding tasks.
 1 Introduction
-Binary code plays an irreplaceable role in the security domain, being the foundation of crucial tasks including vulnerability detection Güler et al.
-(2019); Duan et al.
-(2020); Chen et al.
-(2022b), malware detection Spensky et al.
-(2016); Aonzo et al.
-(2023); Xu et al.
-(2014), binary recovery Su et al.
-(2024); Zhang et al.
-(2021); Chen et al.
-(2022c), and legacy software maintenance Carbone et al.
-(2009); Carlini et al.
-(2015); Martin et al.
-(2010).
-For example, when performing tasks such as identifying attacks and malware, security analysts often only have access to assembly, i.e., the human-readable representation of binary code, which is extremely difficult to understand Su et al.
-(2024); Zhang et al.
-(2021); Chen et al.
-(2022c).
-Thus, combined with the increasing sophistication of cybercrime that poses significant threats worldwide (e.g., cybercrime is predicted to cost the world \$10.5 trillion annually by 2025 Sausalito (2020)), effective binary analysis techniques are in high demand.
+Binary code plays an irreplaceable role in the security domain, being the foundation of crucial tasks including vulnerability detection (Güler et al., 2019; Duan et al., 2020; Chen et al., 2022b), malware detection (Spensky et al., 2016; Aonzo et al., 2023; Xu et al., 2014), binary recovery (Su et al., 2024; Zhang et al., 2021; Chen et al., 2022c), and legacy software maintenance (Carbone et al., 2009; Carlini et al., 2015; Martin et al., 2010).
+For example, when performing tasks such as identifying attacks and malware, security analysts often only have access to assembly, i.e., the human-readable representation of binary code, which is extremely difficult to understand (Su et al., 2024; Zhang et al., 2021; Chen et al., 2022c).
+Thus, combined with the increasing sophistication of cybercrime that poses significant threats worldwide (e.g., cybercrime is predicted to cost the world \$10.5 trillion annually by 2025 (Sausalito, 2020)), effective binary analysis techniques are in high demand.
 Figure 1: Example that shows the semantics and diverse optimizations of assembly code.
-Large language models pre-trained on source code have brought improvement in various software development domains Chen et al.
-(2022a); Liu et al.
-(2023); Chen et al.
-(2023); Le et al.
-(2022); Jiang et al.
-(2023); Xia et al.
-(2023).
-However, these LLMs are not designed for or trained with assembly corpus, not achieving their full potential on binary code analysis tasks such as binary code similarity Wang et al.
-(2022); Xu et al.
-(2023a), malware detection Su et al.
-(2024), and binary code decompilation Tan et al.
-(2024); Armengol-Estapé et al.
-(2024); Hosseini and Dolan-Gavitt (2022).
-Existing work applying LLMs on assembly code mainly piggybacks on encoder-style LLMs Wang et al.
-(2022); Su et al.
-(2024); Xu et al.
-(2023a), unable to benefit from the more extensive pre-training, updated architectures, scaling of state-of-the-art generative LLMs.
-Other work using generative LLMs for decompilation shows a low unit test passing rate of the decompiled programs Tan et al.
-(2024); Armengol-Estapé et al.
-(2024).
+Large language models pre-trained on source code have brought improvement in various software development domains (Chen et al., 2022a; Liu et al., 2023; Chen et al., 2023; Le et al., 2022; Jiang et al., 2023; Xia et al., 2023).
+However, these LLMs are not designed for or trained with assembly corpus, not achieving their full potential on binary code analysis tasks such as binary code similarity (Wang et al., 2022; Xu et al., 2023a), malware detection (Su et al., 2024), and binary code decompilation (Tan et al., 2024; Armengol-Estapé et al., 2024; Hosseini & Dolan-Gavitt, 2022).
+Existing work applying LLMs on assembly code mainly piggybacks on encoder-style LLMs (Wang et al., 2022; Su et al., 2024; Xu et al., 2023a), unable to benefit from the more extensive pre-training, updated architectures, scaling of state-of-the-art generative LLMs.
+Other work using generative LLMs for decompilation shows a low unit test passing rate of the decompiled programs (Tan et al., 2024; Armengol-Estapé et al., 2024).
 The challenges of leveraging generative LLMs for assembly code are twofold.
 First, compared to source code, assembly code has a lower information density.
-A short source-code sequence maps to a much longer assembly-code sequence that is often several times longer.
+A short source-code sequence maps to an assembly-code sequence that is often several times longer.
 Thus, assembly semantics span across a long sequence of tokens.
@@
   • We develop Nova, a generative foundation LLM with hierarchical attention and contrastive learning for assembly.
-Nova outperforms state-of-the-art (SOTA) on binary decompilation by up to 146.54% and on binary similarity detection by up to 6.17%.
-  • We conduct a comprehensive analysis, illustrating the effectiveness of Nova’s novel designs: (1) Nova’s embeddings of assemblies successfully reflect code functionalities in the latent space, and (2) Nova’s hierarchical attention complements standard attention by learning different attention weight distributions, especially those reflecting long sequence semantics.
+Nova outperforms state-of-the-art on binary decompilation by up to 14.84 – 21.58% higher Pass@k and on binary similarity detection by up to 6.17% Recall@1.
 2 Approach
 Figure 2 presents the overall approach of Nova.
-We build Nova on top of foundation models for source code Rozière et al.
-(2023); Li et al.
-(2023); Guo et al.
-(2024) to utilize their source code and natural language generation ability.
+We build Nova on top of foundation models for source code (Guo et al., 2024) to utilize their source code and natural language generation ability.
 We first collect large assembly corpora (Section 2.1).
-Section 2.2 describes Nova’s hierarchical attention.
+Section 2.2 describes Nova’s hierarchical attention design.
 With the collected assembly corpora, we then pretrain Nova with language modeling and contrastive learning objectives (Section 2.3).
@@
 Figure 2: Overview of developing Nova
+Table 1: Statistics (number of source code and assembly functions) of the pre-training datasets.
+Datasets Source O0 O1 O2 O3 Total
+AnghaBench 757.1K 743.1K 726.4K 718.7K 717.8K 3.7M
+The-Stack 138.8K 125.1K 119.7K 116.9K 108.8K 609.3K
 2.1 Data Collection
-We build our assembly data sets on top of existing source code datasets: The-Stack Li et al.
-(2023) and the AnghaBench da Silva et al.
-(2021).
+We build our assembly data sets on top of existing source code datasets: The-Stack (Li et al., 2023) and the AnghaBench (da Silva et al., 2021).
 We compile the source code into executables with different optimization levels (i.e., O0, O1, O2 and O3), strip the executables to remove debug information, and disassemble them into assembly code.
@@
 The breakdown statistics are in Table 2.
-Datasets Source O0 O1 O2 O3 Total
-AnghaBench 757.1K 743.1K 726.4K 718.7K 717.8K 3.7M
-The-Stack 138.8K 125.1K 119.7K 116.9K 108.8K 609.3K
-Table 1: Statistics (number of source code and assembly functions) of the pre-training datasets.
 We perform certain normalization on the assembly functions: (1) removing all the “%” and comments, (2) adding whitespace around “,”, “(”, “)”, (3) converting all the hexadecimal numbers to decimal numbers, and (4) replacing the address of each instruction with special labels (e.g., replacing “0” and “4” in Figure 1 (b) with “[INST-1]” and “[INST-2]”) placing at the end of each instruction.
@@
 Specifically, Nova learns the assembly code in an hierarchical way by providing a modified attention mask.
-Different from standard token-level attentions Vaswani et al.
-(2023); Radford and Narasimhan (2018); Radford et al.
-(2019); Brown et al.
-(2020), our hierarchical self-attention contains three categories at different levels of granularity.
-Figure 3: Design of hierarchical attention
-(1) Intra-Instruction Attention: Due to the low information density in assembly, intra-instruction attention is designed to capture the summary of every instruction, which is the standard causal attention but limited to tokens of each instruction (the yellow part in Figure 3).
+Different from standard token-level attentions (Vaswani et al., 2023; Radford & Narasimhan, 2018; Radford et al., 2019; Brown et al., 2020), our hierarchical self-attention contains three categories at different levels of granularity.
+Figure 3: Design of Nova’s hierarchical attention for assembly code
+(1) Intra-Instruction Attention: Due to the low information density in assembly, intra-instruction attention is designed to capture the summary of every instruction, which is the standard causal attention but limited to tokens of each instruction (the yellow part in Figures 3) (a) and (b).
 Tokens in different instructions have no attention weights.
 The “[INST]” label at the end of the instruction has attention to all the tokens in the instruction and thus captures the semantics of the entire instruction (e.g., “[INST-1]” captures the semantics of “mov eax, \$1”).
-(2) Preceding-Instruction Attention In addition to the local semantics of each instruction, the use of assembly instructions (such as the choice of registers) depends on the context.
+(2) Preceding-Instruction Attention: In addition to the local semantics of each instruction, the use of assembly instructions (such as the choice of registers) depends on the context.
 For example, after the first instruction “mov eax, \$1”, the second instruction should not reuse “eax” to store another value “\$2” immediately.
-To capture such context, the preceding-instruction attention enables each token in an instruction to have attention to the “[INST]” label of the preceding instruction (the light green part in Figure 3).
-(3) Inter-Instruction Attention To understand function semantics (i.e., functionality), which lies in the dependencies across different instructions, the inter-instruction attention is designed to let the “[INST]” label of each instruction have attention to all the labels of previous instructions.
-For example, “[INST-4]” has attention to “[INST-1]”, “[INST-2]”, and “[INST-3]” (the dark green part in Figure 3).
+To capture such context, the preceding-instruction attention enables each token in an instruction to have attention to the “[INST]” label of the preceding instruction (the light green part in Figures 3 (a) and (b)).
+(3) Inter-Instruction Attention: To understand function semantics (i.e., functionality), which lies in the dependencies across different instructions, the inter-instruction attention is designed to let the “[INST]” label of each instruction have attention to all the labels of previous instructions.
+For example, “[INST-4]” has attention to “[INST-1]”, “[INST-2]”, and “[INST-3]” (the dark green part in Figures 3 (a) and (b)).
 The inter-instruction attention is only enabled for “[INST]” labels, as they represent the semantics of each instruction.
@@
 The inter-instruction attention learns the long dependencies across instructions on top of the “[INST]” labels that contain the instruction summary.
-Appendix A.2 shows how hierarchical self-attention works with text and source code.
-Figure 4: Design of functionality and optimization contrastive learning (CL). “asm” denotes assembly.
+Figure 3 (c) also shows how Nova’s hierarchical attention for assembly code is compatible with the standard self-attention for text and source code.
+As existing LLMs have shown good performance on text and source code using the standard self-attention, we keep the standard causal attention mask within and between any chunks of text or source code in the input (the light grey part shown in Figure 3 (c)).
+The attention from text or source code to assembly code (and vice versa) is only allowed for the “[INST]” tokens as they are designed to contain the assembly instruction summaries.
 2.3 Contrastive Learning
 The syntax gap between assembly code and source code, and syntax diversity between differently-optimized assembly code make LLMs struggle to distinguish the semantics behind the syntax.
-Nova adopts contrastive learning technique Gao et al.
-(2021) during pre-training to train LLMs to encode assembly code meaningfully w.r.t semantics.
-The standard pre-training objective is language modeling by minimizing the negative likelihood of code in the pre-training corpus Radford and Narasimhan (2018), notated as $L_{lm}$.
+Nova adopts contrastive learning technique (Gao et al., 2021) during pre-training to train LLMs to encode assembly code meaningfully w.r.t semantics.
+The standard pre-training objective is language modeling by minimizing the negative likelihood of code in the pre-training corpus (Radford & Narasimhan, 2018), notated as $L_{lm}$.
 In addition, Nova is pre-trained with two new objectives, $L_{fcl}$ for functionality contrastive learning and $L_{ocl}$ for optimization contrastive learning.
-Functionality CL Functionality CL trains Nova to focus more on the functionalities of assembly code rather than the syntax.
+Figure 4: Design of functionality and optimization contrastive learning (CL). “asm” denotes assembly.
+Functionality CL: Functionality CL trains Nova to focus more on the functionalities of assembly code rather than the syntax.
 Code with the same functionality (assemblies from the same source code), should be encoded closer in the latent space.
@@
 Once Nova encodes the batch of source code and assembly functions, we calculate the distance matrix $\{D_{ij}\}_{f_{i},f_{j}\in F}=\{d(e^{s}_{f_{i}},e^{t}_{f_{j}})\}$, and minimize the loss:
-$L_{fcl}=-\mathop{log}\sum\limits_{s,t\in S}\sum\limits_{f_{i}\in F}\left(1-\frac{\mathop{exp}\left(d(e^{s}_{f_{i}},e^{t}_{f_{i}})\right)}{\sum_{f_{j}\in F}\mathop{exp}\left(d(e^{s}_{f_{i}},e^{t}_{f_{j}})\right)}\right)$
+$L_{fcl}=-\mathop{log}\sum_{s,t\in S}\sum_{f_{i}\in F}\left(1-\frac{\mathop{exp}\left(d(e^{s}_{f_{i}},e^{t}_{f_{i}})\right)}{\sum_{f_{j}\in F}\mathop{exp}\left(d(e^{s}_{f_{i}},e^{t}_{f_{j}})\right)}\right)$
 This objective minimizes the distance between embeddings for the same function, i.e., the diagonal in the distance matrix.
-Optimization CL LLMs can be confused if being asked to directly connect a source code function to its O3-optimized assembly, due to their dramatically different syntax.
+Optimization CL: LLMs can be confused if being asked to directly connect a source code function to its O3-optimized assembly, due to their dramatically different syntax.
 Such a huge gap can be filled by learning how the source code is transformed to O0, O1, O2 and eventually to O3 assembly, as the optimization levels are ordered.
@@
 Nova learns such order with the optimization CL objective, encoding differently-optimized assembly code orderly.
-Optimization CL optimizes Nova with the constraint:
-$\forall f\in F,\forall s<t_{1}<t_{2}\in S,\;d(e^{s}_{f},e^{t_{1}}_{f})\leq d(e^{s}_{f},e^{t_{2}}_{f})$
-Intuitively, this ensures that the more optimizations applied, the larger the difference between embeddings of optimized and unoptimized code.
+Optimization CL optimizes Nova with the constraint: the more optimizations applied, the larger the difference between embeddings of optimized and unoptimized code.
 For instance, Figure 4 (c) and (d) illustrate that for the same function “cmp”, the distance between source code and assembly increases when the optimization level increases.
@@
 2.5 Task 2: Binary Code Similarity Detection
-Binary code similarity detection (BCSD) aims to measure the similarity between two binary code snippets Wang et al.
-(2022); Su et al.
-(2024), which is the foundation of various applications such as plagiarism detection Luo et al.
-(2014); Sæbjørnsen et al.
-(2009) and vulnerability detection David and Yahav (2014); David et al.
-(2018, 2017, 2016).
+Binary code similarity detection (BCSD) aims to measure the similarity between two binary code snippets (Wang et al., 2022; Su et al., 2024), which is the foundation of various applications such as plagiarism detection (Luo et al., 2014; Sæbjørnsen et al., 2009) and vulnerability detection (David & Yahav, 2014; David et al., 2018; 2017; 2016).
 A widely used setting is taking a query assembly of the function $f^{q}$ that is compiled with one optimization level (denoted by $s$), and a pool of candidate assembly of $K$ functions (notated by $f^{p}_{i},$ $1\leq i\leq K$) compiled with a different optimization level (denoted by $t\neq s$).
@@
 Nova is fine-tuned to encode these binaries, so that the positive candidate has the highest similarity with the query assembly among the pool.
-The learning objective is:
+The learning objective is as follows:
 […]
-, where we follow previous work Su et al. (2024) to let $s$ be O0-assembly and $t$ be O3-assembly, which is the hardest setting.
+We follow previous work (Su et al., 2024) to let $s$ be O0-assembly and $t$ be O3-assembly, which is the hardest setting.
 3 Experimental Setup
+This section describes the setup of pre-training and fine-tuning of Nova, as well as the existing baselines we compare Nova with, and the evaluation metrics we used in the two downstream tasks.
+Appendix A.2 contains additional details such as training hyper-parameters.
 3.1 Pre-Training
 We use the data collected from AnghaBench and The-Stack for pre-training.
-We pre-train Nova starting from DeepSeek-Coder Guo et al.
-(2024), and the hierarchical attention is applied on half of the attention heads to balance between its effectiveness and the existing knowledge in the standard attention layers.
+We pre-train Nova starting from DeepSeek-Coder (Guo et al., 2024), and the hierarchical attention is applied on half of the attention heads to balance between its effectiveness and the existing knowledge in the standard attention layers (Justified in Appendix A.3).
 Nova is pre-trained with language modeling for one epoch, followed by contrastive learning objectives for another epoch.
-3.2 Binary Code Decompilation
-Training Data We sample 2.16M assembly-to-source-code pairs (0.338B tokens) from the pre-training corpus to build the BCD fine-tuning data.
-Test Data We use HumanEval-Decompile Tan et al.
-(2024) as the test benchmark, which was not used in training.
-HumanEval-Decompile is derived from the C language adaptation of the HumanEval Chen et al.
-(2021) benchmark and provides test cases in evaluating functionality correctness.
-HumanEval-Decompile contains 164 C functions, each compiled with O0O3 optimization flags and disassembled into X86-64 assembly.
-Baselines Nova is compared with GPT-3.5, GPT-4, and the previous SOTA LLM4Decompile Tan et al.
-(2024).
-LLM4Decompile trains DeepSeek-Coder using the same AnghaBench corpus, and it is the first LLM-based technique that aims to generate executable decompilations.
-Evaluation GPT-3.5 and GPT-4 are prompted with three-shot examples, while LLM4Decompile and Nova samples 20 decompilations per assembly, using the temperature of 0.2 and top_p of 0.95 Chen et al.
-(2021).
-The generated decompilations are executed against the test cases and both Pass@1 and Pass@10 Chen et al.
-(2021) are reported.
-3.3 Binary Code Similarity Detection
-Training Data To compare Nova with existing works on BCSD fairly Wang et al.
-(2022); Su et al.
-(2024), we use BinaryCorp-3M Wang et al.
-(2022) as the fine-tuning data for BCSD, which contains the O0 and O3 assembly of 224,606 functions.
-Test Data Following existing work Su et al.
-(2024); Xu et al.
-(2023a), we use real-world benchmarks, Binutils, Curl, ImageMagick, SQLite, OpenSSL, and Putty, as the test benchmarks, which are nonexistent in the training data.
-Baselines Nova is compared with jTrans Wang et al.
-(2022), DiEmph Xu et al.
-(2023a) and CodeArt Su et al.
-(2024).
-jTrans is a Transformer Vaswani et al.
-(2023) encoder trained on binaries with masked token prediction and jump target prediction tasks.
+3.2 Fine-Tuning for Binary Code Decompilation
+Training Data: We sample (due to computation resource limitation) 2.16M assembly-to-source-code pairs (0.338B tokens) from the pre-training corpus to build the BCD fine-tuning data.
+Test Data: We use HumanEval-Decompile (Tan et al., 2024) as the test benchmark, which was not used in training.
+HumanEval-Decompile is derived from the C language adaptation of the HumanEval (Chen et al., 2021) benchmark and contains 164 C functions, each compiled with O0O3 optimization flags and disassembled into X86-64 assembly.
+Baselines: Nova is compared with existing SOTA LLM4Decompile (Tan et al., 2024), Meta LLMCompiler (Cummins et al., 2024), and other general code LLMs (CodeLlama (Rozière et al., 2023), StarCoder2 (Lozhkov et al., 2024), DeepSeekCoder (Guo et al., 2024), Qwen2.5-Coder (Hui et al., 2024), Granite Mishra et al.
+(2024), LongCoder (Guo et al., 2023), GPT-3.5-Turbo, and GPT-4o).
+LLM4Decompile trains DeepSeekCoder using the same AnghaBench corpus for binary decompilation.
+Meta LLMCompiler trains CodeLlama models using LLVM IRs, X86, and ARM assembly code to optimize and translate binary code.
+Evaluation: GPT-3.5-Turbo and GPT-4o are prompted with three-shot examples, while all the other LLMs and Nova sample 20 decompilations per assembly code, using the temperature of 0.2 and top_p of 0.95 (Chen et al., 2021).
+The generated decompilations are executed against the test cases.
+Finally, Pass@1 and Pass@10 (Chen et al., 2021) are reported.
+3.3 Fine-Tuning for Binary Code Similarity Detection
+Training Data: To compare Nova with existing works on BCSD fairly (Wang et al., 2022; Su et al., 2024), we use the same dataset, BinaryCorp-3M (Wang et al., 2022), as the fine-tuning data for BCSD, which contains the O0 and O3 assembly of 224,606 functions.
+Test Data: Following existing work (Su et al., 2024; Xu et al., 2023a), we use real-world benchmarks, Binutils, Curl, ImageMagick, SQLite, OpenSSL, and Putty, as the test benchmarks, which are nonexistent in the training data.
+Baselines: Nova is compared with jTrans (Wang et al., 2022), DiEmph (Xu et al., 2023a) and CodeArt (Su et al., 2024).
+jTrans is a Transformer (Vaswani et al., 2023) encoder trained on binaries with masked token prediction and jump target prediction tasks.
 DiEmph uses an instruction deemphasis technique to prevent the model from learning instruction distribution biases introduced by compilers.
 CodeArt proposes a regularized attention mask for encoder models to capture instructional semantics and data dependencies.
-Evaluation We randomly sample $K$ source code functions from each project, compile them into binaries with O0 and O3 optimization flags, and disassemble them into X86-64 assemblies.
-BCSD techniques encode these assemblies into embeddings (Nova uses the average embeddings of all the “[INST]” tokens in an assembly as its embedding).
+Evaluation: We randomly sample $K$ source code functions from each project, compile them into binaries with O0 and O3 optimization flags, and disassemble them into X86-64 assemblies.
+BCSD techniques encode these assemblies into embeddings (Nova uses the average last-layer hidden states of all the “[INST]” tokens in an assembly as its embedding).
 Then each O0 assembly is used as the query to calculate their similarity with the $K$ O3 candidate assemblies (using cosine similarity).
 Metric Recall@1 is the ratio of queries for which the candidate from the same source code has the highest similarity among all the candidates.
-Appendix A.3 contains additional details such as training hyper-parameters.
 4 Results
 4.1 Binary Code Decompilation
-4.1.1 Comparison with SOTA Techniques
-Table 2 shows the Pass@1 of the decompiled code from assemblies on the HumanEval-Decompile benchmark.
-The results are grouped by optimization level (i.e., the benchmark contains 164 assemblies of each optimization level to decompile), and the average is also reported.
-Optimization O0 O1 O2 O3 Avg.
-GPT-3.5 6.80 5.64 4.36 3.93 5.18
-GPT-4 17.77 15.12 12.65 11.25 14.20
-LLM4Decompile-1B 12.26 7.22 8.38 7.96 8.96
-Nova-1B 31.19 17.29 18.72 15.58 22.09
-LLM4Decompile-6B 23.01 15.95 16.95 14.79 17.68
-Nova-6B 42.07 28.04 25.00 22.56 29.42
-Table 2: Pass@1 on HumanEval-Decompile.
-Optimization O0 O1 O2 O3 Avg.
-GPT-3.5 8.95 7.77 5.93 5.12 6.94
-GPT-4 25.64 20.65 18.70 18.03 20.76
-LLM4Decompile-1B 17.95 12.05 13.90 12.51 14.10
-Nova-1B 41.11 29.81 31.18 26.24 32.09
-LLM4Decompile-6B 33.77 24.25 23.94 23.81 26.44
-Nova-6B 51.06 37.83 35.79 34.63 39.83
-Table 3: Pass@10 on HumanEval-Decompile.
-Overall, Nova’s Pass@1 is higher than all SOTA binary decompilation techniques and general LLMs GPT-4 and GPT-3.5, which are orders of magnitude larger than Nova.
-Specifically, for each optimization level, Nova consistently decompiles more assemblies into source code correctly than LLM4Decompile, GPT-3.5, and GPT-4.
-With the same model size, Nova-1B outperforms LLM4Decompile-1B by 146.54%, i.e., averaged Pass@1 of 22.09% versus 8.96%.
-Nova-6B outperforms LLM4Decompile-6B by 66.40%: the averaged Pass@1 is 29.42% versus 17.68%.
-Table 3 shows that Nova still outperforms SOTA techniques with a significant margin under the measurement of Pass@10. Examples of Nova’s correct decompilation are provided in Appendix A.4.
-4.1.2 Ablation Study
-We conduct an ablation study by comparing Nova-1B with the following models:
-  • Nova-CL-HA: Removing contrastive learning and hierarchical self-attention. This is simply training DeepSeek-Coder-1.3B on the assembly corpus using language modeling.
-  • Nova-HA: Removing the hierarchical self-attention, training DeepSeek-Coder-1.3B on the assembly corpus using both the language modeling and contrastive learning objectives.
-Nova-CL-HA can be viewed as our reproduction (retrain) of LLM4Decompile-1B.
-Optimization O0 O1 O2 O3 Avg.
-LLM4Decompile-1B 17.95 12.05 13.90 12.51 14.10
-Nova-CL-HA 17.80 13.32 13.26 10.03 13.60
-Nova-HA 25.12 15.64 16.07 12.71 17.39
-Nova 31.19 17.29 18.72 15.58 22.09
-Table 4: Ablation study of Nova-1B (Pass@1).
-Table 4 shows the results of the ablation study, reported by the Pass@1 on HumanEval-Decompile.
-Nova-CL-HA shows comparable Pass@1, which we considered as variance in reproducing the same approach.
-With additional contrastive learning objectives, Nova-HA improves the Pass@1 on all optimization levels over Nova-CL-HA, showing a 27.87% higher averaged Pass@1.
-Further applying the hierarchical self-attention boosts the overall Pass@1 from 17.39% to 22.09%.
+Comparison with SOTA Techniques: Table 2 shows the Pass@1 and Pass@10 of the decompiled code from assemblies on HumanEval-Decompile.
+The results are grouped by optimization level (i.e., the benchmark contains 164 assemblies of each optimization level), and the average is also reported.
+Table 2: Nova’s Pass@K and comparison with existing techniques on HumanEval-Decompile.
+Techniques Pass@1 Pass@10
+O0 O1 O2 O3 Avg. O0 O1 O2 O3 Avg.
+CodeLlama-7B 6.95 3.81 4.54 3.78 4.77 8.53 5.97 7.34 5.17 6.75
+StarCoder2-7B 6.31 4.33 5.64 5.95 5.56 8.77 5.18 6.09 7.17 6.80
+DeepSeekCoder-7B 9.63 7.56 7.41 6.68 7.82 13.60 11.38 11.52 9.44 11.49
+Qwen-2.5-Coder-7B 4.76 5.79 5.58 5.27 5.35 6.34 7.69 6.56 5.79 6.60
+LLMCompiler-7B 5.95 5.85 5.55 5.82 5.79 7.01 7.31 7.47 7.01 7.20
+GPT-3.5-Turbo 7.41 6.13 4.33 3.90 5.44 9.56 8.38 6.23 5.12 7.32
+GPT-4o 21.34 18.29 14.48 13.05 16.79 29.94 26.74 21.42 19.88 24.50
+LLM4Decompile-1B 15.30 8.26 9.36 8.38 10.33 21.79 15.23 16.17 13.70 16.72
+Nova-1B 37.53 21.71 22.68 18.75 25.17 49.38 34.84 36.95 32.03 38.30
+LLM4Decompile-6B 29.97 19.05 20.46 18.32 21.95 40.40 27.75 28.85 28.51 31.38
+Nova-6B 48.78 30.58 30.85 27.23 34.36 57.47 47.45 43.03 39.68 46.91
+Overall, Nova’s Pass@1 and Pass@10 are higher than all SOTA binary decompilation techniques and general LLMs with even smaller model sizes.
+Specifically, for each optimization level, Nova consistently decompiles more assemblies into source code correctly than the rest of the compared techniques.
+Note that Meta LLMCompiler is mainly designed for LLVM IR optimization, and thus is still incapable of assembly code decompilation.
+With the same model size, Nova-1B outperforms LLM4Decompile-1B with a 14.84% higher averaged Pass@1, and a 21.58% higher Pass@10.
+Nova-6B outperforms LLM4Decompile-6B with a 12.41% higher averaged Pass@1, and a 15.53% higher Pass@10.
+When compared with GPT-4o, an order of magnitude larger model, Nova-1B produces an 8.38% higher Pass@1 and 13.80% higher Pass@10.
+Examples of Nova’s correct decompilation are provided in Appendix A.4.
+Comparison with Techniques Handling Long Input: Nova’s hierarchical attention design targets to address the low information density and long input challenge of assembly code.
+There are other techniques that handles long input challenges in text and source code, with Granite-3B-Code-Base-128K and LongCoder being the most related ones.
+Granite trains LLM on repository-level long inputs, which is an orthogonal approach with Nova’s approach (hierarchical attention and contrastive learning).
+We train Granite-3B-Code-128K with Nova’s approach, and Table 3 shows that Nova’s approach brings improvement to Granite over standard fine-tuning even if it has already been trained with long code data.
+Table 3: Nova’s approach brings improvement to LLM that has been trained with long input data.
+Techniques Pass@1 Pass@10
+O0 O1 O2 O3 Avg. O0 O1 O2 O3 Avg.
+Granite (3B-Code-128K) 5.91 3.78 5.09 5.52 5.08 8.19 5.16 6.56 7.15 6.76
+Granite + Standard Fine-Tuning 20.88 13.54 11.37 10.09 13.97 30.05 19.77 18.31 15.77 20.98
+Granite + Nova’s Fine-Tuning 31.04 14.57 14.70 13.66 18.49 39.57 21.23 21.77 19.82 25.60
+Table 4: Nova’s hierarchical attention is more effective on assembly code.
+Techniques Pass@1 Pass@10
+O0 O1 O2 O3 Avg. O0 O1 O2 O3 Avg.
+DeepSeekCoder + LongCoder’s Attention 34.59 19.07 19.72 17.34 22.68 42.19 32.37 32.86 29.04 34.12
+DeepSeekCoder + Nova’s Attention 37.53 21.71 22.68 18.75 25.17 49.38 34.84 36.95 32.03 38.30
+LongCoder combines window attention and global attention to learn long code input.
+We compare LongCoder’s attention design with Nova’s hierarchical attention design using DeepSeekCoder-1.3B as the backbone.
+Table 4 shows that Nova’s hierarchical attention is more effective in learning assembly code.
+Nova’s attention design considers the instruction-level local semantics and dependencies between different instructions, which fits better than fix-sized window attention to assembly code.
+Ablation Study: We conduct an ablation study by comparing Nova-1B with the following models to show the effectiveness of contrastive learning objectives and hierarchical attention design:
+  • Nova-CL-HA: Removing contrastive learning and hierarchical self-attention.
+This is simply training DeepSeekCoder-1.3B on the assembly corpus using language modeling.
+This can be viewed as our reproduction (retrain) of LLM4Decompile-1B using the same amount of data.
+  • Nova-HA: Removing the hierarchical self-attention, training DeepSeekCoder-1.3B on the assembly corpus using both the language modeling and contrastive learning objectives.
+Table 5: Ablation study of Nova-1B on HumanEval-Decompile.
+Techniques Pass@1 Pass@10
+O0 O1 O2 O3 Avg. O0 O1 O2 O3 Avg.
+LLM4Decompile-1B 15.30 8.26 9.36 8.38 10.33 21.79 15.23 16.17 13.70 16.72
+Nova-CL-HA 20.73 16.16 15.03 11.19 15.78 33.55 28.12 26.96 21.01 27.41
+Nova-HA 30.58 19.88 20.58 16.40 21.86 44.75 33.13 33.31 29.82 35.25
+Nova 37.53 21.71 22.68 18.75 25.17 49.38 34.84 36.95 32.03 38.30
+Table 5 shows the results of the ablation study.
+Nova-CL-HA produces an average Pass@1 of 15.78% and Pass@10 of 27.41%.
+With additional contrastive learning objectives, Nova-HA improves the Pass@1 on all optimization levels over Nova-CL-HA, showing a higher averaged Pass@1 and Pass@10.
+Further applying the hierarchical self-attention on Nova-HA boosts the overall Pass@1 from 21.86% to 25.17%, and Pass@10 from 35.25% to 38.30%.
+Table 6: Recall@1 on BCSD with $K=50$
 Benchmarks jTrans DiEmph CodeArt Nova-1B Nova-6B
@@
 Avg. 0.67 0.78 0.81 0.84 0.86
-Table 5: Recall@1 on benchmarks with $K=50$.
+Table 7: Recall@1 on BCSD with $K=100$
 Benchmarks jTrans DiEmph CodeArt Nova-1B Nova-6B
@@
 Avg. 0.60 0.72 0.77 0.78 0.79
-Table 6: Recall@1 on benchmarks with $K=100$.
+Table 8: Recall@1 on BCSD with $K=200$
 Benchmarks jTrans DiEmph CodeArt Nova-1B Nova-6B
@@
 Avg. 0.51 0.64 0.71 0.73 0.75
-Table 7: Recall@1 on benchmarks with $K=200$.
+Table 9: Recall@1 on BCSD with $K=500$
 Benchmarks jTrans DiEmph CodeArt Nova-1B Nova-6B
@@
 Avg. 0.39 0.55 0.64 0.64 0.67
-Table 8: Recall@1 on benchmarks with $K=500$.
 4.2 Binary Code Similarity Detection
-4.2.1 Comparison with SOTA Techniques
-Tables 8,  8,  8 and  8 show the Recall@1 of Nova and existing BCSD techniques with pool size $K$ of 50, 100, 200 and 500 on the six benchmarks.
-Underlined numbers indicates the best in each benchmark, while wavey underlined numbers denote the tied best (we only mark Nova-1B for clearer illustration).
-Overall, Tables 888 and 8 show that on average, Nova-1B and Nova-6B achieve the highest Recall@1 (in bold) under all four settings of $K$.
+Tables 9,  9,  9 and  9 show the Recall@1 of Nova and existing BCSD techniques with pool size $K$ of 50, 100, 200 and 500 on the six benchmarks.
+Underline indicates the best in each benchmark, while wave denotes the tied best (we only mark Nova-1B for clearer illustration).
+Overall, Tables 9,  9,  9 and 9 show that on average, Nova-1B and Nova-6B achieve the highest Recall@1 (in bold) under all four settings of $K$.
 Nova-6B further outperforms Nova-1B and achieves the highest averaged Recall@1 under all four settings, ranking the ground-truth of 5%, 2%, 4%, and 3% more queries the most similar correspondingly compared to CodeArt.
@@
 For instance, Nova-1B wins on four benchmarks while DiEmph only wins on SQLite when $K=50$.
-$K$ Nova-CL-HA Nova-HA Nova
-50 0.81 0.83 (+0.02) 0.84 (+0.01)
-100 0.76 0.78 (+0.02) 0.78
-200 0.70 0.70 0.73 (+0.03)
-500 0.60 0.62 (+0.02) 0.64 (+0.02)
-Table 9: Ablation study of Nova-1B (Recall@1)
-4.2.2 Ablation Study
-Table 9 shows the averaged Recall@1 of Nova-CL-HA, Nova-HA (same as in Section 4.1.2), and Nova-1B under four pool size settings.
-With contrastive learning objectives, Nova-HA improves Nova-CL-HA under three settings ($K=50,100,200$) with 2% higher Recall@1.
-With hierarchical attention, Nova further outperforms Nova-HA under three settings ($K=50,200,500$).
-Detailed ablation study results on each benchmark are provided in Appendix A.5.
-4.3 Analytic Experiments
-4.3.1 How are Nova’s embeddings better?
+We also conduct an ablation study on BCSD in Appendix A.5.
+4.3 Analytic Experiments: How are Nova’s embeddings better?
 […]
-Figure 5 (b) shows that Nova encodes assemblies into clusters of functionalities.
-The assemblies for the same functionality (i.e., the same task) are encoded closer to each other than Nova-CL-HA does in Figure 5 (b).
-The results show that our hierarchical attention and contrastive learning techniques effectively group codes of similar functionalities together for better assembly foundation models.
-Embedding of Nova-HA and additional quantitative analysis are shown in Appendix A.6.
-Figure 5: PCA of embeddings calculated by Nova-CL-HA and Nova, for HumanEval-Decompile assemblies.
-4.3.2 What does hierarchical attention learn?
-We conduct quantitative analysis on the attention weights produced by different models.
-Figure 6: Quantitative analysis of attention weights.
-Entropy Figure 6 (a) shows the entropy of attention-weight distributions in each layer.
-We separate the attention heads as standard attention (green line) and hierarchical attention (red line), since Nova applies hierarchical attention to half of the attention heads in each layer (Section 3.1).
-Nova’s hierarchical attention heads produce significantly lower entropy, suggesting its attention layer is more confident in learning specific relationships than the other models’ attention layers.
-The standard attention heads in Nova show patterns similar to those of Nova-CL-HA and Nova-HA, allowing Nova learning the standard attention to capture the “soft” relationship between possible tokens pairs.
-The result suggests that hierarchical attention complements standard attention with additional knowledge.
-[INST] Token Figure 6 (b) shows the attention weights paid to the “[INST]” tokens.
-Nova’s hierarchical attention heads pay more attention to the “[INST]” tokens than standard attention, which may be because these “[INST]” tokens contain instruction summary and long dependencies and thus are more informative.
-Additional analysis and examples are given in Appendix A.7.
+Compared with Nova-CL-HA (Figure 5 (a)), Nova-HA (Figure 5 (b)) including contrastive learning objectives in the pre-training, can separate the embeddings of assemblies with different functionalities better.
+Nova-HA clearly encode “Task 143” (orange points) away from the others.
+Yet, Nova’s (Figure 5 (c)) embeddings group the assemblies by functionalities more precisely than Nova-HA, suggesting that hierarchical attention enhances the training of contrastive learning objectives to learn more effective encoding.
+Visualization using a different approach, PCA, is shown in Appendix A.6.
+Analytic experiments on Nova’s hierarchical attention is shown in Appendix A.7.
+Figure 5: t-SNE analysis of embeddings cauclated by Nova-CL-HA, Nova-HA, and Nova.
 5 Related Work
@@
 Machine learning models are widely used in binary program analysis tasks.
-However, these models are typically designed for specific downstream tasks such as binary code similarity detection Pei et al.
-(2020); Xu et al.
-(2023a); Wang et al.
-(2022), variable name prediction Chen et al.
-(2022c); Xu et al.
-(2023b); Zhang et al.
-(2021), binary code type inference Pei et al.
-(2021), and are built from scratch.
+However, these models are typically designed for specific tasks such as binary code similarity detection (Pei et al., 2020; Xu et al., 2023a; Wang et al., 2022), variable name prediction (Chen et al., 2022c; Xu et al., 2023b; Zhang et al., 2021), binary code type inference (Pei et al., 2021), and are built from scratch.
 Recent techniques have started to pre-train foundation LLMs for binaries.
-CodeArt Su et al.
-(2024) pre-trains encoder-style LLMs with a regularized attention design to better encode assembly code semantics, showing good accuracy on binary code understanding tasks (e.g., binary code similarity detection and malware family classification).
-SLaDe Armengol-Estapé et al.
-(2024) trains BART Lewis et al.
-(2019) models on assembly, and LLM4Decompile Tan et al.
-(2024) trains DeepSeek-Coder with assembly for binary code decompilation.
+CodeArt (Su et al., 2024) pre-trains encoder-style LLMs with a regularized attention design to better encode assembly code semantics.
+SLaDe (Armengol-Estapé et al., 2024) trains BART (Lewis et al., 2019) models on assembly.
+Meta LLMCompiler (Cummins et al., 2024) train CodeLlama models on LLVM IR to optimize binary code.
+LLM4Decompile (Tan et al., 2024) trains DeepSeekCoder with assembly for binary code decompilation.
 However, CodeArt does not generalize to generation tasks due to its encoder architecture.
+LLMCompiler trained on LLVM IR cannot be effectively transferred to assembly code.
 SLaDe and LLM4Decompile are limited in performance due to a lack of special designs for assembly.
-In contrast, Nova addresses both limitations, by using the proposed hierarchical attention and contrastive learning objectives, outperforming existing techniques on both understanding (binary code similarity detection) and generation (binary code decompilation) tasks.
+In contrast, Nova proposes hierarchical attention and contrastive learning objectives, outperforming existing techniques on both understanding (binary code similarity detection) and generation (binary code decompilation) tasks.
 5.2 Large Source-Code Models
-LLMs demonstrate promising results on many code-related tasks, such as code generation Chen et al.
-(2022a); Liu et al.
-(2023); Chen et al.
-(2023); Le et al.
-(2022); Yue et al.
-(2021); Chen et al.
-(2021); Nijkamp et al.
-(2022); Fried et al.
-(2023); Rozière et al.
-(2023); Guo et al.
-(2024), bug fixing Jiang et al.
-(2023); Xia et al.
-(2023) and vulnerability fixing Wu et al.
-(2023); Steenhoek et al.
-(2023); He and Vechev (2023).
-The advances in using LLMs are attributed to the knowledge learned from massive source code and natural language text in their training datasets Touvron et al.
-(2023); OpenAI (2023).
+LLMs demonstrate promising results on many code-related tasks, such as code generation (Chen et al., 2022a; Liu et al., 2023; Chen et al., 2023; Le et al., 2022; Yue et al., 2021; Chen et al., 2021; Nijkamp et al., 2022; Fried et al., 2023; Rozière et al., 2023; Guo et al., 2024; Lozhkov et al., 2024; Hui et al., 2024), bug fixing (Jiang et al., 2023; Xia et al., 2023) and vulnerability fixing (Wu et al., 2023; Steenhoek et al., 2023; He & Vechev, 2023).
+The advances in using LLMs are attributed to the knowledge learned from massive source code and natural language text in their training datasets (Touvron et al., 2023; OpenAI, 2023).
 Nova is designed and trained for assembly, which has unique challenges such as low information density and diverse optimization.
-6 Conclusion
-This work develops Nova, a generative foundation LLM for assembly code, which incorporates two key novelties (hierarchical attention and contrastive learning objectives) to address the unique challenges of assembly code.
-Evaluation on downstream tasks shows the effectiveness of Nova, which outperforms existing techniques on binary code decompilation by up to 146.54% and outperforms the latest binary code similarity detection techniques by up to 6.17%.
-We expect our hierarchical attention and contrastive learning techniques to benefit source code and natural language foundation models, which remains as future work.
+5.3 Attention Mechanism
+Standard self-attention is widely used in transformer architecture (Vaswani et al., 2023) to capture soft dependencies between tokens in the input.
+Many special attention mechanisms have been designed for better learning in various scenarios (Yang et al., 2016; Huang et al., 2024).
+LongCoder (Guo et al., 2023) combines window attention and global attention (attention sink (Xiao et al., 2024)) to handle long input of source code.
+We have shown that LongCoder’s window attention is less effective than Nova’s on assembly code.
+CAST (Shi et al., 2021) is a new neural architecture that splits the abstract syntax tree (AST) of source code into subtrees, encodes the subtrees, and aggregates to the final encoding.
+PA-former (Chai & Li, 2024) is a new neural architecture that constructs source code as pyramid input based on their AST structure and contains a pyramid attention mechanism to calculate the features in a hierarchical way.
+HierarchyNet (Nguyen et al., 2023) is a neural architecture that considers source code AST, data flow, and control flow graphs.
+Similarly, it cannot be applied to assembly code.
+Different from CAST, PA-former, and HierarchiyNet, Nova’s attention design is for assembly code, is more lightweight and can be plugged into any pre-trained generative LLM.
 6 Limitations
 One limitation is that Nova is X86-specific, as we only collect X86 assembly corpus for pre-training.
-This design choice is mainly affected by two facts: (1) X86 assembly is used and explored in a wide range of binary tasks Wang et al.
-(2022); Su et al.
-(2024); Xu et al.
-(2023a); Chen et al.
-(2022c) compared to other assembly languages, and (2) computation limitations.
+This design choice is mainly affected by two facts: (1) X86 assembly is used and explored in a wide range of binary tasks (Wang et al., 2022; Su et al., 2024; Xu et al., 2023a; Chen et al., 2022c) compared to other assembly languages, and (2) computation limitations.
 However, the proposed techniques are independent of X86 assembly.
@@
 However, due to the computing resources limitation, we are unable to explore that in this work.
+7 Conclusion
+This work develops Nova, a generative foundation LLM for assembly code, which incorporates two key novelties (hierarchical attention and contrastive learning objectives) to address the unique challenges of assembly code.
+Evaluation on downstream tasks shows the effectiveness of Nova, which outperforms existing techniques on binary code decompilation by up to 146.54% and outperforms the latest binary code similarity detection techniques by up to 6.17%.
+We expect our hierarchical attention and contrastive learning techniques to benefit source code and natural language foundation models, which remains as future work.
 ACKNOWLEDGMENTS
 This research was supported in part by NSF 1901242 and 2006688 and a CFI fund.
+This paper was prepared for informational purposes in part by the CDAO group of JPMorgan Chase & Co and its affiliates (”J.P.
+Morgan”) and is not a product of the Research Department of J.P.
+Morgan.
+J.P.
+Morgan makes no representation and warranty whatsoever and disclaims all liability, for the completeness, accuracy or reliability of the information contained herein.
+This document is not intended as investment research or investment advice, or a recommendation, offer or solicitation for the purchase or sale of any security, financial instrument, financial product or service, or to be used in any way for evaluating the merits of participating in any transaction, and shall not constitute a solicitation under any jurisdiction or to any person, if such solicitation under such jurisdiction or to such person would be unlawful.
 Appendix A Appendix
@@
 Figure 6: Example of assembly code preprocessing
-A.2 Hierarchical Self-Attention
-The hierarchical self-attention is designed for assembly code, yet the input to LLMs may still contain text or source code.
-Figure 8 illustrates how the hierarchical attention works with text or source code in the input.
-As existing LLMs have shown good performance on text and source code using the standard self-attention, we keep the standard causal attention mask within and between any chunks of text or source code in the input (the light grey part shown in Figure 8).
-The attention from text or source code to assembly code (and vice versa) is only allowed for the “[INST]” tokens as they are supposed to contain the assembly instruction summaries.
-Figure 8: Hierarchical attention with text input.
+A.2 Training Details
+This section provides additional details of training.
+We pre-train Nova starting from DeepSeek-Coder, using the language modeling objective ($L_{lm}$) for one epoch on the AnghaBench and The-Stack corpora.
+This step uses a batch size of 128, with the input truncated by a 1,024 tokens limit.
+The model weights are updated using the AdamW optimizer.
+The learning rate is $5e^{-5}$, using 1000 steps of warm-up and a cosine decay to adjust the learning rate.
+Then, the model is further pre-trained with the combination of language modeling and contrastive learning objectives ($L=L_{lm}+\lambda(L_{fcl}+L_{ocl})$), with $\lambda$ set to 0.1.
+To train with the functionality contrastive learning objective, we discard any source code that misses any one of O0O3 assemblies and also discard the source code whose O2 assembly is the same as its O3 assembly.
+As a result, this step is only trained for 0.36M data samples for one epoch.
+The batch size is 64, with the input truncated by a 1,024 tokens limit.
+The learning rate is $2e^{-5}$ using the AdamW optimizer.
+The fine-tuning of both BCD and BCSD uses a batch size of 64, with the input truncated by a 2,048 token limit.
+Similarly, the learning rate is $2e^{-5}$ using the AdamW optimizer, and the model is fine-tuned for one epoch.
+During the training using the contrastive learning objectives, and the fine-tuning of BCSD, we use the average of [INST] tokens’ last layer hidden states to represent the embedding of a binary function.
+Infrastructure The training are conducted on eight NVIDIA RTX A100 GPUs, each with 40GB memory.
+Our implementation is based on Huggingface’s implementation of DeepSeek-Coder  [^1] https://huggingface.co/deepseek-ai/deepseek-coder-1.3b-base, PyTorch  [^2] https://pytorch.org/get-started/locally/, and DeepSpeed  [^3] https://github.com/microsoft/DeepSpeed.
+Table 10: Comparison with applying hierarchical attention on all attention heads, using 1B models.
+Techniques Pass@1 Pass@10
+O0 O1 O2 O3 Avg. O0 O1 O2 O3 Avg.
+Nova-HA 30.58 19.88 20.58 16.40 21.86 44.75 33.13 33.31 29.82 35.25
+Nova (hierarchical on all heads) 32.38 18.87 20.56 16.34 22.04 45.95 32.19 32.78 29.01 34.98
+Nova 37.53 21.71 22.68 18.75 25.17 49.38 34.84 36.95 32.03 38.30
+A.3 Applying Hierarchical Attention on Half Attention Heads
+The hierarchical attention mask is always applied on half of the attention heads at each layer in Nova.
+This ensures the LLM balances the hierarchical knowledge of assembly code and pre-trained knowledge learned by full self-attention.
+Figure 7: Nova-1B correctly decompiles HumanEval-Decompile task 0.
+We conducted experiments applying hierarchical attention to all the attention heads.
+Results in Table 10 show that when applying hierarchical attention to all the attention heads of transformer layers, the performance does not improve enough and even drops under some settings.
+This implies that the standard full self-attention mechanism indeed learns knowledge that may not be captured by hierarchical attention.
+Thus, we only apply hierarchical attention to half of the attention heads in each transformer layer of Nova to balance the knowledge learned by standard and hierarchical attention.
+Figure 8: Nova-1B correctly decompiles HumanEval-Decompile task 20.
 Benchmarks Nova-CL-HA Nova-HA Nova-1B
@@
 Table 14: Ablation study with $K=500$.
-A.3 Training Details
-This section provides additional details of training.
-We pre-train Nova starting from DeepSeek-Coder, using the language modeling objective ($L_{lm}$) for one epoch on the AnghaBench and The-Stack corpora.
-This step uses a batch size of 128, with the input truncated by a 1,024 tokens limit.
-The model weights are updated using the AdamW optimizer.
-The learning rate is $5e^{-5}$, using 1000 steps of warm-up and a cosine decay to adjust the learning rate.
-Then, the model is further pre-trained with the combination of language modeling and contrastive learning objectives ($L=L_{lm}+\lambda(L_{fcl}+L_{ocl})$), with $\lambda$ set to 0.1.
-To train with the functionality contrastive learning objective, we discard any source code that misses any one of O0O3 assemblies and also discard the source code whose O2 assembly is the same as its O3 assembly.
-As a result, this step is only trained for 0.36M data samples for one epoch.
-The batch size is 64, with the input truncated by a 1,024 tokens limit.
-The learning rate is $2e^{-5}$ using the AdamW optimizer.
-The fine-tuning of both BCD and BCSD uses a batch size of 64, with the input truncated by a 2,048 token limit.
-Similarly, the learning rate is $2e^{-5}$ using the AdamW optimizer, and the model is fine-tuned for one epoch.
-Infrastructure The training are conducted on eight NVIDIA RTX A100 GPUs, each with 40GB memory.
-Our implementation is based on Huggingface’s implementation of DeepSeek-Coder  [^1] https://huggingface.co/deepseek-ai/deepseek-coder-1.3b-base, PyTorch  [^2] https://pytorch.org/get-started/locally/, and DeepSpeed  [^3] https://github.com/microsoft/DeepSpeed.
-Figure 9: Nova-1B correctly decompiles HumanEval-Decompile task 0, while GPT-4 and LLM4Decompile-1B fail.
-Figure 10: Nova-1B correctly decompiles HumanEval-Decompile task 20, while LLM4Decompile-1B fail.
-Figure 11: PCA of embeddings calculated by Nova-CL-HA, Nova-HA, and Nova.
-Figure 12: t-SNE of embeddings cauclated by Nova-CL-HA, Nova-HA, and Nova.
 A.4 Binary Code Decompilation Case Studies
@@
 Nova-1B correctly decompiles the source code, successfully figuring that the function is trying to find the two elements that are closest to each other in the given array *a, with the result stored in minmax.
+Figure 9: PCA of embeddings calculated by Nova-CL-HA, Nova-HA, and Nova.
+Figure 10: Comparison of attention distribution among standard and hierarchical heads.
 A.5 Binary Code Similarity Detection Ablation Study
-Table $13$ show the detailed ablation study results of BCSD. Nova wins on the most benchmarks when $K=100$ or $500$, and ties with Nova-HA when $K=50$, or $200$.
+Table $14$ show the detailed ablation study results of BCSD. Nova wins on the most benchmarks when $K=100$ or $500$, and ties with Nova-HA when $K=50$, or $200$.
 A.6 Additional Analysis of Embedding
-Additional Embedding Visualization
-Figure 11 shows the full results of PCA of embeddings provided by Nova-CL-HA, Nova-HA, and Nova, on randomly sampled seven examples.
-Compared with Nova-CL-HA, Nova-HA including contrastive learning objectives in the pre-training, can separate the embeddings of assemblies with different functionalities better.
-Nova-HA clearly encode “Task 143” (orange points) away from the others.
-Nova’s embeddings group the assemblies by functionalities more precisely than Nova-HA, suggesting that hierarchical attention enhances the training of contrastive learning objectives to learn more effective encoding.
-Figure 12 shows the results using another dimensionality reduction technique, t-SNE van der Maaten and Hinton (2008), where Nova’s embeddings are consistently more distinguishable by functionalities.
-Figure 13: Distances between embeddings of different-formats-same-functionality code and same-format-different-functionalities code.
-Figure 14: Comparison of attention distribution among standard and hierarchical heads in the final layer.
-Quantitative Analysis
-We also perform quantitative analysis to further support the conclusion from visualization.
-Figure 13 (a) shows the distribution of $l_{2}$ distances between different-formats-same-functionality code in HumanEval-Decompile (“format” is defined as one of source code, O0, O1, O2, or O3 assembly).
-The figure shows five distributions, (src, O0), (src, O1), (src, O2), and (src, O3).
-Each distribution calculates the distances between embeddings of two formats of code, e.g., src, O0 refers to the source code and O0 assembly for the same task in HumanEval-Decompile.
-Figure 13 (a) shows that Nova’s embeddings for same-functionality-different-formats code are closer to each other in the latent space.
-For the same task, all the O0O3 assemblies’ embeddings have a smaller $l_{2}$ distance to the source code embedding compared to that of Nova-CL-HA’s and Nova-HA’s embeddings.
-Another interesting finding is that for Nova, the distances between assembly and source code increase with the optimization level of assemblies increases.
-This trend matches what the optimization CL tries to optimize for.
-Figure 13 (b) shows the distribution of $l_{2}$ distance between same-format-different-functionalities code from the HumanEval-Decompile benchmark.
-With the same format (source code or the same optimized assembly), Nova encodes the assemblies with different functionalities farther away from each other compared to Nova-CL-HA.
-Nova-CL-HA’s embeddings cannot significantly reflect the functionality differences between assemblies, while Nova’s embeddings can.
-Nova-HA’s embeddings show high $l_{2}$ distance in both Figure 13 (a) and (b), suggesting that Nova-HA tries to decentralize all the code in the embedding space even if they have the same functionality, which is not as desired as Nova.
+Figure 9 shows the results of PCA of embeddings provided by Nova-CL-HA, Nova-HA, and Nova, on randomly sampled seven examples, where Nova’s embeddings are consistently more distinguishable by functionalities.
 A.7 Additional Analysis of Attention
@@
 In addition, in lower layers, we have observed attention weights to be softly distributed among tokens comprising each instruction (Figure 11), which suggests Nova initially models cross-relations among operation codes and operands in the first few layers, and later pools their summary representation into the [INST] token in the later layers.
-This is also supported by Figure 6, where the Nova’s hierarchical attention (red line) shows a decreasing trend of entropy.
-This means the hierarchical attention is softer in lower layers, and becomes focused in higher layers.

### v6

@@
 Figure 1: Example that shows the semantics and diverse optimizations of assembly code.
-Large language models pre-trained on source code have brought improvement in various software development domains (Chen et al., 2022a; Liu et al., 2023; Chen et al., 2023; Le et al., 2022; Jiang et al., 2023; Xia et al., 2023).
+Large language models pre-trained on source code have brought improvement in various software development domains (Chen et al., 2022a; Liu et al., 2023a; Chen et al., 2023; Le et al., 2022; Jiang et al., 2023; Xia et al., 2023).
 However, these LLMs are not designed for or trained with assembly corpus, not achieving their full potential on binary code analysis tasks such as binary code similarity (Wang et al., 2022; Xu et al., 2023a), malware detection (Su et al., 2024), and binary code decompilation (Tan et al., 2024; Armengol-Estapé et al., 2024; Hosseini & Dolan-Gavitt, 2022).
@@
   • We develop Nova, a generative foundation LLM with hierarchical attention and contrastive learning for assembly.
-Nova outperforms state-of-the-art on binary decompilation by up to 14.84 – 21.58% higher Pass@k and on binary similarity detection by up to 6.17% Recall@1.
+Nova outperforms state-of-the-art on binary decompilation by up to 14.84 – 21.58% higher Pass@1 and Pass@10, and on binary similarity detection by up to 6.17% Recall@1.
+2 Related Work
+2.1 Binary Models
+Machine learning models are widely used in binary program analysis tasks.
+However, these models are typically designed for specific tasks such as binary code similarity detection (Pei et al., 2020; Xu et al., 2023a; Wang et al., 2022; Xu et al., 2017; Ding et al., 2019), variable name prediction (Chen et al., 2022c; Xu et al., 2023b; Zhang et al., 2021; He et al., 2018; Lacomis et al., 2020), binary code type inference (Pei et al., 2021), and so on (Chen et al., 2022d; Liu et al., 2023b; Hosseini & Dolan-Gavitt, 2022).
+Recent techniques have started to pre-train foundation LLMs for binaries.
+CodeArt (Su et al., 2024) pre-trains encoder-style LLMs with a regularized attention design to better encode assembly code semantics.
+SLaDe (Armengol-Estapé et al., 2024) trains BART (Lewis et al., 2019) models on assembly.
+Meta LLMCompiler (Cummins et al., 2024) train CodeLlama models on LLVM IR to optimize binary code.
+LLM4Decompile (Tan et al., 2024) trains DeepSeekCoder with assembly for binary code decompilation.
+However, CodeArt does not generalize to generation tasks due to its encoder architecture.
+LLMCompiler trained on LLVM IR cannot be effectively transferred to assembly code.
+SLaDe and LLM4Decompile are limited in performance due to a lack of special designs for assembly.
+In contrast, Nova proposes hierarchical attention and contrastive learning objectives, outperforming existing techniques on both understanding (binary code similarity detection) and generation (binary code decompilation) tasks.
+2.2 Large Source-Code Models
+LLMs demonstrate promising results on many code-related tasks, such as code generation (Chen et al., 2022a; Liu et al., 2023a; Chen et al., 2023; Le et al., 2022; Yue et al., 2021; Chen et al., 2021; Nijkamp et al., 2022; Fried et al., 2023; Rozière et al., 2023; Guo et al., 2024; Lozhkov et al., 2024; Hui et al., 2024), bug fixing (Jiang et al., 2023; Xia et al., 2023) and vulnerability fixing (Wu et al., 2023; Steenhoek et al., 2023; He & Vechev, 2023).
+The advances in using LLMs are attributed to the knowledge learned from massive source code and natural language text in their training datasets (Touvron et al., 2023; OpenAI, 2023).
+Nova is designed and trained for assembly, which has unique challenges such as low information density and diverse optimization.
+2.3 Attention Mechanism
+Standard self-attention is widely used in transformer architecture (Vaswani et al., 2017) to capture soft dependencies between tokens in the input.
+Many special attention mechanisms have been designed for better learning in various scenarios (Yang et al., 2016; Huang et al., 2024).
+LongCoder (Guo et al., 2023) combines window attention and global attention (attention sink (Xiao et al., 2024)) to handle long input of source code.
+We have shown that LongCoder’s window attention is less effective than Nova’s on assembly code.
+CAST (Shi et al., 2021) is a new neural architecture that splits the abstract syntax tree (AST) of source code into subtrees, encodes the subtrees, and aggregates to the final encoding.
+PA-former (Chai & Li, 2024) is a new neural architecture that constructs source code as pyramid input based on their AST structure and contains a pyramid attention mechanism to calculate the features in a hierarchical way.
+HierarchyNet (Nguyen et al., 2023) is a neural architecture that considers source code AST, data flow, and control flow graphs.
+Similarly, it cannot be applied to assembly code.
+Different from CAST, PA-former, and HierarchiyNet, Nova’s attention design is for assembly code, is more lightweight and can be plugged into any pre-trained generative LLM.
 3 Approach
@@
 With the collected assembly corpora, we then pretrain Nova with language modeling and contrastive learning objectives (Section 3.3).
-Then, we fine-tune Nova on two important downstream tasks, binary code decompilation, and binary code similarity detection (Sections 2.4 and 2.5), to prove Nova’s effectiveness and benefits to the binary research domain.
+Then, we fine-tune Nova on two important downstream tasks, binary code decompilation, and binary code similarity detection (Sections 3.4 and 3.5), to prove Nova’s effectiveness and benefits to the binary research domain.
 Figure 2: Overview of developing Nova
-Table 1: Statistics (number of source code and assembly functions) of the pre-training datasets.
-Datasets Source O0 O1 O2 O3 Total
+Table 1: Statistics (number of C and X86-64 assembly functions) of the pre-training datasets.
+Origin C Functions O0 O1 O2 O3 Total
 AnghaBench 757.1K 743.1K 726.4K 718.7K 717.8K 3.7M
@@
 3.1 Data Collection
-We build our assembly data sets on top of existing source code datasets: The-Stack (Li et al., 2023) and the AnghaBench (da Silva et al., 2021).
-We compile the source code into executables with different optimization levels (i.e., O0, O1, O2 and O3), strip the executables to remove debug information, and disassemble them into assembly code.
-We treat every assembly function as a separate data sample.
-The breakdown statistics are in Table 2.
+In this paper, we focus on X86-64 assembly functions for C programs. Yet, Nova’s approach is generalizable to other assembly languages such as ARM assembly.
+We derive our X86-64 assembly functions dataset from two source code corpora: C functions in The-Stack (Li et al., 2023) and AnghaBench (da Silva et al., 2021).
+We compile the C programs into executables using gcc with different optimization levels (i.e., O0, O1, O2 and O3), strip the executables to remove debug information, and disassemble them into X86-64 assembly code using objdump.
+We treat every function as a separate data point.
+Table 3 shows the number of C functions in the two original datasets, and the number of X86-64 assembly functions we collected from them.
 We perform certain normalization on the assembly functions: (1) removing all the “%” and comments, (2) adding whitespace around “,”, “(”, “)”, (3) converting all the hexadecimal numbers to decimal numbers, and (4) replacing the address of each instruction with special labels (e.g., replacing “0” and “4” in Figure 1 (b) with “[INST-1]” and “[INST-2]”) placing at the end of each instruction.
@@
 Specifically, Nova learns the assembly code in an hierarchical way by providing a modified attention mask.
-Different from standard token-level attentions (Vaswani et al., 2023; Radford & Narasimhan, 2018; Radford et al., 2019; Brown et al., 2020), our hierarchical self-attention contains three categories at different levels of granularity.
+Different from standard token-level attentions (Vaswani et al., 2017; Radford & Narasimhan, 2018; Radford et al., 2019; Brown et al., 2020), our hierarchical self-attention contains three categories at different levels of granularity.
 Figure 3: Design of Nova’s hierarchical attention for assembly code
@@
 The inter-instruction attention is only enabled for “[INST]” labels, as they represent the semantics of each instruction.
-To sum up, the hierarchical self-attention breaks the semantics of assembly code into three parts.
-The intra-instruction attention captures the instruction summary, and the preceding-instruction attention captures the context with the preceding instruction.
-The inter-instruction attention learns the long dependencies across instructions on top of the “[INST]” labels that contain the instruction summary.
-Figure 3 (c) also shows how Nova’s hierarchical attention for assembly code is compatible with the standard self-attention for text and source code.
-As existing LLMs have shown good performance on text and source code using the standard self-attention, we keep the standard causal attention mask within and between any chunks of text or source code in the input (the light grey part shown in Figure 3 (c)).
-The attention from text or source code to assembly code (and vice versa) is only allowed for the “[INST]” tokens as they are designed to contain the assembly instruction summaries.
+To sum up, the hierarchical self-attention splits assembly code semantics into three levels: intra-instruction attention captures instruction summaries, preceding-instruction attention provides context from the preceding instruction, and inter-instruction attention models long dependencies across instructions using “[INST]” tokens that contain the instruction summary.
+Figure 3 (c) highlights the compatibility of Nova’s hierarchical attention with standard self-attention for text and source code.
+Leveraging the proven performance of standard self-attention in existing LLMs, we retain the causal attention mask within and across chunks of text or source code (shown in light grey in Figure 3 (c)).
+Cross-attention between text, source code, and assembly is restricted to “[INST]” tokens, which encapsulate assembly instruction summaries.
 3.3 Contrastive Learning
@@
 For instance, in Figure 4 (a), embeddings of source and assembly code of function “cmp” are closer to each other, and the same for function “sort”.
+Nova is designed and implemented on decoder-only generative LLMs, and we refer the hidden states from the last transformer layer as embedding.
+For source code, we use the average of each token’s embedding as the source code function’s embedding.
+For assembly, we use the average of all the “[INST]” tokens’ embedding as the embedding of the assembly function, as each “[INST]” token is supposed to capture the semantics of that instruction by the design of our hierarchical self-attention.
 Let $e^{s}_{f}$ be the embedding of function $f$ in $s$ form ($s=-1$ for source code, and $s\in[0,1,2,3]$ for O0 to O3 optimized assembly).
 For simplicity, let $S=[-1,0,1,2,3]$ be the domain of $s$.
-We use the average of all the “[INST]” tokens’ embedding as the embedding of the whole assembly function, as each “[INST]” token is supposed to capture the semantics of that instruction by the design of hierarchical self-attention.
-Functionality CL optimizes Nova with the constraint:
+Functionality CL optimizes Nova’s embeddings to satisfy the constraint:
 […]
 […]
-Such constraints can be trained by optimizing the embeddings of a batch of functions, each function in two different forms.
+The embeddings of a batch of functions, each represented in two different forms, can be optimized to satisfy these constraints.
 For the example in Figure 4 (b), there are two forms (source code and O0 assembly) of four functions.
@@
 $L_{fcl}=-\mathop{log}\sum_{s,t\in S}\sum_{f_{i}\in F}\left(1-\frac{\mathop{exp}\left(d(e^{s}_{f_{i}},e^{t}_{f_{i}})\right)}{\sum_{f_{j}\in F}\mathop{exp}\left(d(e^{s}_{f_{i}},e^{t}_{f_{j}})\right)}\right)$
-This objective minimizes the distance between embeddings for the same function, i.e., the diagonal in the distance matrix.
+This objective minimizes the distance between the embeddings of the same function, which is the diagonal in the distance matrix.
+In practice, this loss is computed and back-propagated for each batch of functions.
+Given a batch size of 64, $F$ represents a set of 64 unique functions in the batch.
 Optimization CL: LLMs can be confused if being asked to directly connect a source code function to its O3-optimized assembly, due to their dramatically different syntax.
@@
 Binary code similarity detection (BCSD) aims to measure the similarity between two binary code snippets (Wang et al., 2022; Su et al., 2024), which is the foundation of various applications such as plagiarism detection (Luo et al., 2014; Sæbjørnsen et al., 2009) and vulnerability detection (David & Yahav, 2014; David et al., 2018; 2017; 2016).
-A widely used setting is taking a query assembly of the function $f^{q}$ that is compiled with one optimization level (denoted by $s$), and a pool of candidate assembly of $K$ functions (notated by $f^{p}_{i},$ $1\leq i\leq K$) compiled with a different optimization level (denoted by $t\neq s$).
+A widely used setting is taking a query assembly of the function $f^{q}$ that is compiled with one optimization level (denoted by $s$), and a pool of candidate assembly of $K$ (e.g., 50, 100, etc.) functions (notated by $f^{p}_{i},$ $1\leq i\leq K$) compiled with a different optimization level (denoted by $t\neq s$).
 There exists a unique candidate assembly coming from the same source code as the query ($\exists!1\leq i\leq K,$ $f^{p}_{i}=f^{q}$, called the positive candidate).
@@
 This section describes the setup of pre-training and fine-tuning of Nova, as well as the existing baselines we compare Nova with, and the evaluation metrics we used in the two downstream tasks.
-Appendix A.2 contains additional details such as training hyper-parameters.
+Appendix A.2 contains additional details such as training hyper-parameters, and evaluation setup.
 4.1 Pre-Training
-We use the data collected from AnghaBench and The-Stack for pre-training.
+We use the C and X86-64 assembly functions collected from AnghaBench and The-Stack for pre-training.
 We pre-train Nova starting from DeepSeek-Coder (Guo et al., 2024), and the hierarchical attention is applied on half of the attention heads to balance between its effectiveness and the existing knowledge in the standard attention layers (Justified in Appendix A.3).
@@
 HumanEval-Decompile is derived from the C language adaptation of the HumanEval (Chen et al., 2021) benchmark and contains 164 C functions, each compiled with O0O3 optimization flags and disassembled into X86-64 assembly.
-Baselines: Nova is compared with existing SOTA LLM4Decompile (Tan et al., 2024), Meta LLMCompiler (Cummins et al., 2024), and other general code LLMs (CodeLlama (Rozière et al., 2023), StarCoder2 (Lozhkov et al., 2024), DeepSeekCoder (Guo et al., 2024), Qwen2.5-Coder (Hui et al., 2024), Granite Mishra et al.
-(2024), LongCoder (Guo et al., 2023), GPT-3.5-Turbo, and GPT-4o).
+Baselines: Nova is compared with existing SOTA LLM4Decompile (Tan et al., 2024), Meta LLMCompiler (Cummins et al., 2024), other open-sourced general code LLMs (Rozière et al., 2023; Lozhkov et al., 2024; Guo et al., 2024; Hui et al., 2024; Mishra et al., 2024; Guo et al., 2023), and commercial LLMs GPT-3.5-Turbo and GPT-4o).
 LLM4Decompile trains DeepSeekCoder using the same AnghaBench corpus for binary decompilation.
 Meta LLMCompiler trains CodeLlama models using LLVM IRs, X86, and ARM assembly code to optimize and translate binary code.
-Evaluation: GPT-3.5-Turbo and GPT-4o are prompted with three-shot examples, while all the other LLMs and Nova sample 20 decompilations per assembly code, using the temperature of 0.2 and top_p of 0.95 (Chen et al., 2021).
-The generated decompilations are executed against the test cases.
+Evaluation: We let each model sample 20 decompilations per assembly function, using the temperature of 0.2 and top_p of 0.95 (Chen et al., 2021).
+Except for LLM4Decompile and Nova that are fine-tuned for binary code decompilation, we provide all other baselines with three-shot examples for few-shot learning (Brown et al., 2020).
+The decompilations are executed with the test cases to verify the functional correctness.
 Finally, Pass@1 and Pass@10 (Chen et al., 2021) are reported.
@@
 Baselines: Nova is compared with jTrans (Wang et al., 2022), DiEmph (Xu et al., 2023a) and CodeArt (Su et al., 2024).
-jTrans is a Transformer (Vaswani et al., 2023) encoder trained on binaries with masked token prediction and jump target prediction tasks.
+jTrans is a Transformer (Vaswani et al., 2017) encoder trained on binaries with masked token prediction and jump target prediction tasks.
 DiEmph uses an instruction deemphasis technique to prevent the model from learning instruction distribution biases introduced by compilers.
 CodeArt proposes a regularized attention mask for encoder models to capture instructional semantics and data dependencies.
-Evaluation: We randomly sample $K$ source code functions from each project, compile them into binaries with O0 and O3 optimization flags, and disassemble them into X86-64 assemblies.
+Evaluation: We randomly sample $K=$ 50, 100, 200, 500 source code functions from each project, compile them into binaries with O0 and O3 optimization flags, and disassemble them into X86-64 assemblies.
 BCSD techniques encode these assemblies into embeddings (Nova uses the average last-layer hidden states of all the “[INST]” tokens in an assembly as its embedding).
-Then each O0 assembly is used as the query to calculate their similarity with the $K$ O3 candidate assemblies (using cosine similarity).
-Metric Recall@1 is the ratio of queries for which the candidate from the same source code has the highest similarity among all the candidates.
+Then each O0 assembly is used as the query to calculate their similarity with the $K$ O3 candidate assemblies.
+Metric Recall@1 is reported as the ratio of queries for which the candidate from the same source code has the highest similarity among all the candidates.
 5 Results
@@
 GPT-4o 21.34 18.29 14.48 13.05 16.79 29.94 26.74 21.42 19.88 24.50
-LLM4Decompile-1B 15.30 8.26 9.36 8.38 10.33 21.79 15.23 16.17 13.70 16.72
-Nova-1B 37.53 21.71 22.68 18.75 25.17 49.38 34.84 36.95 32.03 38.30
-LLM4Decompile-6B 29.97 19.05 20.46 18.32 21.95 40.40 27.75 28.85 28.51 31.38
-Nova-6B 48.78 30.58 30.85 27.23 34.36 57.47 47.45 43.03 39.68 46.91
+LLM4Decompile-1.3B 15.30 8.26 9.36 8.38 10.33 21.79 15.23 16.17 13.70 16.72
+Nova-1.3B 37.53 21.71 22.68 18.75 25.17 49.38 34.84 36.95 32.03 38.30
+LLM4Decompile-6.7B 29.97 19.05 20.46 18.32 21.95 40.40 27.75 28.85 28.51 31.38
+Nova-6.7B 48.78 30.58 30.85 27.23 34.36 57.47 47.45 43.03 39.68 46.91
 Overall, Nova’s Pass@1 and Pass@10 are higher than all SOTA binary decompilation techniques and general LLMs with even smaller model sizes.
@@
 Note that Meta LLMCompiler is mainly designed for LLVM IR optimization, and thus is still incapable of assembly code decompilation.
-With the same model size, Nova-1B outperforms LLM4Decompile-1B with a 14.84% higher averaged Pass@1, and a 21.58% higher Pass@10.
-Nova-6B outperforms LLM4Decompile-6B with a 12.41% higher averaged Pass@1, and a 15.53% higher Pass@10.
-When compared with GPT-4o, an order of magnitude larger model, Nova-1B produces an 8.38% higher Pass@1 and 13.80% higher Pass@10.
+With the same model size, Nova-1.3B outperforms LLM4Decompile-1.3B with a 14.84% higher averaged Pass@1, and a 21.58% higher Pass@10.
+Nova-6.7B outperforms LLM4Decompile-6.7B with a 12.41% higher averaged Pass@1, and a 15.53% higher Pass@10.
+When compared with GPT-4o, an order of magnitude larger model, Nova-1.3B produces an 8.38% higher Pass@1 and 13.80% higher Pass@10.
 Examples of Nova’s correct decompilation are provided in Appendix A.5.
@@
 Granite + Standard Fine-Tuning 20.88 13.54 11.37 10.09 13.97 30.05 19.77 18.31 15.77 20.98
-Granite + Nova’s Fine-Tuning 31.04 14.57 14.70 13.66 18.49 39.57 21.23 21.77 19.82 25.60
+Granite + Nova’s Approaches 31.04 14.57 14.70 13.66 18.49 39.57 21.23 21.77 19.82 25.60
 Table 4: Nova’s hierarchical attention is more effective on assembly code.
@@
 O0 O1 O2 O3 Avg. O0 O1 O2 O3 Avg.
-DeepSeekCoder + LongCoder’s Attention 34.59 19.07 19.72 17.34 22.68 42.19 32.37 32.86 29.04 34.12
-DeepSeekCoder + Nova’s Attention 37.53 21.71 22.68 18.75 25.17 49.38 34.84 36.95 32.03 38.30
+Nova-1.3B (using LongCoder’s Attention) 34.59 19.07 19.72 17.34 22.68 42.19 32.37 32.86 29.04 34.12
+Nova-1.3B 37.53 21.71 22.68 18.75 25.17 49.38 34.84 36.95 32.03 38.30
 LongCoder combines window attention and global attention to learn long code input.
-We compare LongCoder’s attention design with Nova’s hierarchical attention design using DeepSeekCoder-1.3B as the backbone.
+We compare LongCoder’s attention design with Nova’s hierarchical attention design by replacing the hierarchical attention of Nova-1.3B with LongCoder’s attention design.
 Table 4 shows that Nova’s hierarchical attention is more effective in learning assembly code.
 Nova’s attention design considers the instruction-level local semantics and dependencies between different instructions, which fits better than fix-sized window attention to assembly code.
-Ablation Study: We conduct an ablation study by comparing Nova-1B with the following models to show the effectiveness of contrastive learning objectives and hierarchical attention design:
+Ablation Study: We conduct an ablation study by comparing Nova-1.3B with the following models to show the effectiveness of contrastive learning objectives and hierarchical attention design:
   • Nova-CL-HA: Removing contrastive learning and hierarchical self-attention.
 This is simply training DeepSeekCoder-1.3B on the assembly corpus using language modeling.
-This can be viewed as our reproduction (retrain) of LLM4Decompile-1B using the same amount of data.
+This can be viewed as our reproduction (retrain) of LLM4Decompile-1.3B using the same amount of data.
   • Nova-HA: Removing the hierarchical self-attention, training DeepSeekCoder-1.3B on the assembly corpus using both the language modeling and contrastive learning objectives.
-Table 5: Ablation study of Nova-1B on HumanEval-Decompile.
+Table 5: Ablation study of Nova-1.3B on HumanEval-Decompile.
 Techniques Pass@1 Pass@10
 O0 O1 O2 O3 Avg. O0 O1 O2 O3 Avg.
-LLM4Decompile-1B 15.30 8.26 9.36 8.38 10.33 21.79 15.23 16.17 13.70 16.72
+LLM4Decompile-1.3B 15.30 8.26 9.36 8.38 10.33 21.79 15.23 16.17 13.70 16.72
 Nova-CL-HA 20.73 16.16 15.03 11.19 15.78 33.55 28.12 26.96 21.01 27.41
@@
 Table 6: Recall@1 on BCSD with $K=50$
-Benchmarks jTrans DiEmph CodeArt Nova-1B Nova-6B
+Benchmarks jTrans DiEmph CodeArt Nova-1.3B Nova-6.7B
 Binutils 0.68 0.80 0.84 0.87 0.89
@@
 Table 7: Recall@1 on BCSD with $K=100$
-Benchmarks jTrans DiEmph CodeArt Nova-1B Nova-6B
+Benchmarks jTrans DiEmph CodeArt Nova-1.3B Nova-6.7B
 Binutils 0.60 0.63 0.81 0.79 0.79
@@
 Table 8: Recall@1 on BCSD with $K=200$
-Benchmarks jTrans DiEmph CodeArt Nova-1B Nova-6B
+Benchmarks jTrans DiEmph CodeArt Nova-1.3B Nova-6.7B
 Binutils 0.51 0.64 0.74 0.73 0.73
@@
 Table 9: Recall@1 on BCSD with $K=500$
-Benchmarks jTrans DiEmph CodeArt Nova-1B Nova-6B
+Benchmarks jTrans DiEmph CodeArt Nova-1.3B Nova-6.7B
 Binutils 0.40 0.57 0.70 0.65 0.67
@@
 Tables 9,  9,  9 and  9 show the Recall@1 of Nova and existing BCSD techniques with pool size $K$ of 50, 100, 200 and 500 on the six benchmarks.
-Underline indicates the best in each benchmark, while wave denotes the tied best (we only mark Nova-1B for clearer illustration).
-Overall, Tables 9,  9,  9 and 9 show that on average, Nova-1B and Nova-6B achieve the highest Recall@1 (in bold) under all four settings of $K$.
-Nova-6B further outperforms Nova-1B and achieves the highest averaged Recall@1 under all four settings, ranking the ground-truth of 5%, 2%, 4%, and 3% more queries the most similar correspondingly compared to CodeArt.
-Nova-1B consistently outperforms existing techniques with higher Recall@1 when $K$ is 50, 100, and 200, meaning it correctly ranks ground-truth of 3%, 1%, and 2% more queries as the most similar.
-Under the setting of $K=500$, Nova-1B ties with CodeArt with the same highest Recall@1.
-When looking into each individual benchmark, Nova-1B always wins on the most benchmarks under different settings of pool size $K$.
-For instance, Nova-1B wins on four benchmarks while DiEmph only wins on SQLite when $K=50$.
+Underline indicates the best in each benchmark, while wave denotes the tied best (we only mark Nova-1.3B for clearer illustration).
+Overall, Tables 9,  9,  9 and 9 show that on average, Nova-1.3B and Nova-6.7B achieve the highest Recall@1 (in bold) under all four settings of $K$.
+Nova-6.7B further outperforms Nova-1.3B and achieves the highest averaged Recall@1 under all four settings, ranking the ground-truth of 5%, 2%, 4%, and 3% more queries the most similar correspondingly compared to CodeArt.
+Nova-1.3B consistently outperforms existing techniques with higher Recall@1 when $K$ is 50, 100, and 200, meaning it correctly ranks ground-truth of 3%, 1%, and 2% more queries as the most similar.
+Under the setting of $K=500$, Nova-1.3B ties with CodeArt with the same highest Recall@1.
+When looking into each individual benchmark, Nova-1.3B always wins on the most benchmarks under different settings of pool size $K$.
+For instance, Nova-1.3B wins on four benchmarks while DiEmph only wins on SQLite when $K=50$.
 We also conduct an ablation study on BCSD in Appendix A.6.
@@
 Analytic experiments on Nova’s hierarchical attention is shown in Appendix A.8.
-Figure 5: t-SNE analysis of embeddings cauclated by Nova-CL-HA, Nova-HA, and Nova.
-5 Related Work
-5.1 Binary Models
-Machine learning models are widely used in binary program analysis tasks.
-However, these models are typically designed for specific tasks such as binary code similarity detection (Pei et al., 2020; Xu et al., 2023a; Wang et al., 2022), variable name prediction (Chen et al., 2022c; Xu et al., 2023b; Zhang et al., 2021), binary code type inference (Pei et al., 2021), and are built from scratch.
-Recent techniques have started to pre-train foundation LLMs for binaries.
-CodeArt (Su et al., 2024) pre-trains encoder-style LLMs with a regularized attention design to better encode assembly code semantics.
-SLaDe (Armengol-Estapé et al., 2024) trains BART (Lewis et al., 2019) models on assembly.
-Meta LLMCompiler (Cummins et al., 2024) train CodeLlama models on LLVM IR to optimize binary code.
-LLM4Decompile (Tan et al., 2024) trains DeepSeekCoder with assembly for binary code decompilation.
-However, CodeArt does not generalize to generation tasks due to its encoder architecture.
-LLMCompiler trained on LLVM IR cannot be effectively transferred to assembly code.
-SLaDe and LLM4Decompile are limited in performance due to a lack of special designs for assembly.
-In contrast, Nova proposes hierarchical attention and contrastive learning objectives, outperforming existing techniques on both understanding (binary code similarity detection) and generation (binary code decompilation) tasks.
-5.2 Large Source-Code Models
-LLMs demonstrate promising results on many code-related tasks, such as code generation (Chen et al., 2022a; Liu et al., 2023; Chen et al., 2023; Le et al., 2022; Yue et al., 2021; Chen et al., 2021; Nijkamp et al., 2022; Fried et al., 2023; Rozière et al., 2023; Guo et al., 2024; Lozhkov et al., 2024; Hui et al., 2024), bug fixing (Jiang et al., 2023; Xia et al., 2023) and vulnerability fixing (Wu et al., 2023; Steenhoek et al., 2023; He & Vechev, 2023).
-The advances in using LLMs are attributed to the knowledge learned from massive source code and natural language text in their training datasets (Touvron et al., 2023; OpenAI, 2023).
-Nova is designed and trained for assembly, which has unique challenges such as low information density and diverse optimization.
-5.3 Attention Mechanism
-Standard self-attention is widely used in transformer architecture (Vaswani et al., 2023) to capture soft dependencies between tokens in the input.
-Many special attention mechanisms have been designed for better learning in various scenarios (Yang et al., 2016; Huang et al., 2024).
-LongCoder (Guo et al., 2023) combines window attention and global attention (attention sink (Xiao et al., 2024)) to handle long input of source code.
-We have shown that LongCoder’s window attention is less effective than Nova’s on assembly code.
-CAST (Shi et al., 2021) is a new neural architecture that splits the abstract syntax tree (AST) of source code into subtrees, encodes the subtrees, and aggregates to the final encoding.
-PA-former (Chai & Li, 2024) is a new neural architecture that constructs source code as pyramid input based on their AST structure and contains a pyramid attention mechanism to calculate the features in a hierarchical way.
-HierarchyNet (Nguyen et al., 2023) is a neural architecture that considers source code AST, data flow, and control flow graphs.
-Similarly, it cannot be applied to assembly code.
-Different from CAST, PA-former, and HierarchiyNet, Nova’s attention design is for assembly code, is more lightweight and can be plugged into any pre-trained generative LLM.
-6 Limitations
-One limitation is that Nova is X86-specific, as we only collect X86 assembly corpus for pre-training.
-This design choice is mainly affected by two facts: (1) X86 assembly is used and explored in a wide range of binary tasks (Wang et al., 2022; Su et al., 2024; Xu et al., 2023a; Chen et al., 2022c) compared to other assembly languages, and (2) computation limitations.
-However, the proposed techniques are independent of X86 assembly.
-Low information density and compiler optimization are the common challenges of most assembly languages such as X86, ARM, and MIPS.
-The proposed techniques can be applied to the future development of multi-lingual assembly LLMs.
-Another potential limitation is the scale of models.
-We develop Nova-1B and Nova-6B.
-These two sized LLMs show impressive ability in assembly code decompilation and encoding.
-There might be potential benefit of developing larger Nova models.
-However, due to the computing resources limitation, we are unable to explore that in this work.
+Figure 5: t-SNE analysis of embeddings calculated by Nova-CL-HA, Nova-HA, and Nova.
 6 Conclusion
-This work develops Nova, a generative foundation LLM for assembly code, which incorporates two key novelties (hierarchical attention and contrastive learning objectives) to address the unique challenges of assembly code.
-Evaluation on downstream tasks shows the effectiveness of Nova, which outperforms existing techniques on binary code decompilation by up to 146.54% and outperforms the latest binary code similarity detection techniques by up to 6.17%.
-We expect our hierarchical attention and contrastive learning techniques to benefit source code and natural language foundation models, which remains as future work.
+This work presents Nova, a generative foundation LLM for assembly code that introduces two key innovations—hierarchical attention and contrastive learning—to tackle the unique challenges of assembly code understanding.
+Nova builds upon a source code LLM and undergoes further pre-training on approximately 4.3 million collected assembly functions.
+We fine-tune and evaluate Nova on two downstream tasks: binary code decompilation and binary code similarity detection.
+Nova achieves up to 14.84 – 21.58% (absolute percentage point improvement) higher Pass@1 and Pass@10 over existing methods in binary code decompilation and up to a 6.17% higher Recall@1 in binary code similarity detection.
+Looking forward, we believe that the proposed hierarchical attention and contrastive learning methods hold promise for enhancing foundation models in both source code and natural language domains, which we leave as future work.
 ACKNOWLEDGMENTS
-This research was supported in part by NSF 1901242 and 2006688 and a CFI fund.
-This paper was prepared for informational purposes in part by the CDAO group of JPMorgan Chase & Co and its affiliates (”J.P.
-Morgan”) and is not a product of the Research Department of J.P.
-Morgan.
-J.P.
-Morgan makes no representation and warranty whatsoever and disclaims all liability, for the completeness, accuracy or reliability of the information contained herein.
-This document is not intended as investment research or investment advice, or a recommendation, offer or solicitation for the purchase or sale of any security, financial instrument, financial product or service, or to be used in any way for evaluating the merits of participating in any transaction, and shall not constitute a solicitation under any jurisdiction or to any person, if such solicitation under such jurisdiction or to such person would be unlawful.
+This research was supported in part by NSF 1901242 and 2006688, a CFI fund, and J.P.
+Morgan AI Faculty Research Awards.
+Any opinions, findings, and conclusions in this paper are those of the authors only and do not necessarily reflect the views of our sponsors.
 Appendix A Appendix
@@
 Thus, the number of O0O1 we obtain from each corpus is different and smaller than the number of source codes as shown in Table 3.
+There may be alternative ways of collecting assembly code, e.g., using gcc -S to directly dump the assembly code without producing executables.
+However, a key difference is that the assembly generated by gcc -S does not undergo the linking step.
+In practice, binary decompilation and analysis are typically performed on executable or linked assembly code, as it includes linker modifications and reflects the final binary structure.
+Our data collection aligns better with the practical use of assembly.
+Nevertheless, Nova’s approaches should be generalizable to assembly code obtained with a different approach, and the experiments remain as future work.
 Figure 6 shows an example of preprocessing the raw assembly code as described in Section 3.1.
 Figure 6: Example of assembly code preprocessing
-A.2 Training Details
+A.2 Experimental Setup Details
 This section provides additional details of training.
-We pre-train Nova starting from DeepSeek-Coder, using the language modeling objective ($L_{lm}$) for one epoch on the AnghaBench and The-Stack corpora.
+We pre-train Nova starting from DeepSeek-Coder, using the language modeling objective ($L_{lm}$) for one epoch on the C functions and assembly functions collected from AnghaBench and The-Stack corpora.
 This step uses a batch size of 128, with the input truncated by a 1,024 tokens limit.
@@
 During the training using the contrastive learning objectives, and the fine-tuning of BCSD, we use the average of [INST] tokens’ last layer hidden states to represent the embedding of a binary function.
-Infrastructure The training are conducted on eight NVIDIA RTX A100 GPUs, each with 40GB memory.
+For the evaluation of BCSD, we reuse the framework provided by CodeArt Su et al.
+(2024) to evaluate the binary code similarity detection results once Nova produces the embeddings for functions in the test dataset.
+For each one of the $K$ functions, the O0 assembly function is used as the query to calculate the cosine similarity between its embedding and the embeddings of all the $K$ O3 assembly functions.
+Note that Nova does not normalize the embeddings of assembly functions during training and uses $l_{2}$ distance to calculate the $L_{BCSD}$, which optimize Nova to embed the O0 and O3 assembly functions from the same source code have the smallest $l_{2}$ distances.
+When using CodeArt’s framework for evaluation which ranks the similarity using the cosine similarity, we normalize Nova’s embedding for each assembly function since $l_{2}$ distance after normalization keeps the same order as cosine similarity (smaller $l_{2}$ distance means higher cosine similarity).
+Infrastructure The experiments are conducted on eight NVIDIA RTX A100 GPUs, each with 40GB memory.
 Our implementation is based on Huggingface’s implementation of DeepSeek-Coder  [^1] https://huggingface.co/deepseek-ai/deepseek-coder-1.3b-base, PyTorch  [^2] https://pytorch.org/get-started/locally/, and DeepSpeed  [^3] https://github.com/microsoft/DeepSpeed.
@@
 This ensures the LLM balances the hierarchical knowledge of assembly code and pre-trained knowledge learned by full self-attention.
-Figure 7: Nova-1B correctly decompiles HumanEval-Decompile task 0.
 We conducted experiments applying hierarchical attention to all the attention heads.
@@
 Thus, we only apply hierarchical attention to half of the attention heads in each transformer layer of Nova to balance the knowledge learned by standard and hierarchical attention.
-Figure 8: Nova-1B correctly decompiles HumanEval-Decompile task 20.
-Benchmarks Nova-CL-HA Nova-HA Nova-1B
+Table 11: Ablation study of Nova-1.3B on HumanEval-Decompile.
+Techniques Pass@1 Pass@10
+O0 O1 O2 O3 Avg. O0 O1 O2 O3 Avg.
+Nova-CL-HA 20.73 16.16 15.03 11.19 15.78 33.55 28.12 26.96 21.01 27.41
+Nova-FCL-HA 22.38 16.20 16.37 13.25 17.05 36.13 29.48 30.02 23.76 29.85
+Nova-OCL-HA 28.44 18.87 18.53 15.76 20.40 40.28 32.33 31.80 27.05 32.87
+Nova-HA 30.58 19.88 20.58 16.40 21.86 44.75 33.13 33.31 29.82 35.25
+Nova 37.53 21.71 22.68 18.75 25.17 49.38 34.84 36.95 32.03 38.30
+A.4 Additional Ablation Study on Binary Code Decompilation
+We provide additional ablation studies on studying the impact of each individual contrastive learning objective. We study two more models:
+  • Nova-FCL-HA: Removing functional contrastive learning and hierarchical self-attention.
+  • Nova-OCL-HA: Removing optimization contrastive learning and hierarchical self-attention.
+Table 11 shows that each component of the contrastive learning brings certain improvements to the Pass@1 and Pass@10 on HumanEval-Decompile, and we find the impact of function contrastive learning (FCL) is larger than the impact of optimization contrastive learning (OCL), suggesting that aligning the model’s embeddings for assembly code with the same functionality is more useful.
+Figure 7: Nova-1.3B correctly decompiles HumanEval-Decompile task 0.
+Figure 8: Nova-1.3B correctly decompiles HumanEval-Decompile task 20.
+A.5 Binary Code Decompilation Case Studies
+Figure 7 shows an example from HumanEval-Decompile (task_id 0).
+Given the O1-optimized assembly code, GPT-4 fails to figure out the number of function arguments correctly, missing one important argument “float e”, and thus produces wrong functionality in the decompiled code.
+LLM4Decompile-1.3B makes similar mistakes and also misses the inner nested for loop.
+Nova-1.3B correctly decompiles the assembly into source code, where the ground truth is checking if any two elements in the given list *a (with size n) are close to each other than a given threshold e.
+Figure 8 shows another more complex example, HumanEval-Decompile task_id 20.
+Nova-1.3B correctly decompiles the source code, successfully figuring that the function is trying to find the two elements that are closest to each other in the given array *a, with the result stored in minmax.
+Table 12: Ablation study with $K=50$.
+Benchmarks Nova-CL-HA Nova-HA Nova-1.3B
 Binutils 0.86 0.88 0.87
@@
 Avg. 0.81 0.83 0.84
-Table 11: Ablation study with $K=50$.
-Benchmarks Nova-CL-HA Nova-HA Nova-1B
+Table 13: Ablation study with $K=100$.
+Benchmarks Nova-CL-HA Nova-HA Nova-1.3B
 Binutils 0.80 0.82 0.79
@@
 Avg. 0.76 0.78 0.78
-Table 12: Ablation study with $K=100$.
-Benchmarks Nova-CL-HA Nova-HA Nova-1B
+Table 14: Ablation study with $K=200$.
+Benchmarks Nova-CL-HA Nova-HA Nova-1.3B
 Binutils 0.71 0.74 0.73
@@
 Avg. 0.70 0.70 0.73
-Table 13: Ablation study with $K=200$.
-Benchmarks Nova-CL-HA Nova-HA Nova-1B
+Table 15: Ablation study with $K=500$.
+Benchmarks Nova-CL-HA Nova-HA Nova-1.3B
 Binutils 0.62 0.65 0.65
@@
 Avg. 0.60 0.62 0.64
-Table 14: Ablation study with $K=500$.
-A.4 Binary Code Decompilation Case Studies
-Figure 7 shows an example from HumanEval-Decompile (task_id 0).
-Given the O1-optimized assembly code, GPT-4 fails to figure out the number of function arguments correctly, missing one important argument “float e”, and thus produces wrong functionality in the decompiled code.
-LLM4Decompile-1B makes similar mistakes and also misses the inner nested for loop.
-Nova-1B correctly decompiles the assembly into source code, where the ground truth is checking if any two elements in the given list *a (with size n) are close to each other than a given threshold e.
-Figure 8 shows another more complex example, HumanEval-Decompile task_id 20.
-Nova-1B correctly decompiles the source code, successfully figuring that the function is trying to find the two elements that are closest to each other in the given array *a, with the result stored in minmax.
+A.6 Binary Code Similarity Detection Ablation Study
+Table $15$ show the detailed ablation study results of BCSD. Nova wins on the most benchmarks when $K=100$ or $500$, and ties with Nova-HA when $K=50$, or $200$.
 Figure 9: PCA of embeddings calculated by Nova-CL-HA, Nova-HA, and Nova.
+A.7 Additional Analysis of Embedding
+Figure 9 shows the results of PCA of embeddings provided by Nova-CL-HA, Nova-HA, and Nova, on randomly sampled seven examples, where Nova’s embeddings are consistently more distinguishable by functionalities.
 Figure 10: Comparison of attention distribution among standard and hierarchical heads.
-A.5 Binary Code Similarity Detection Ablation Study
-Table $14$ show the detailed ablation study results of BCSD. Nova wins on the most benchmarks when $K=100$ or $500$, and ties with Nova-HA when $K=50$, or $200$.
-A.6 Additional Analysis of Embedding
-Figure 9 shows the results of PCA of embeddings provided by Nova-CL-HA, Nova-HA, and Nova, on randomly sampled seven examples, where Nova’s embeddings are consistently more distinguishable by functionalities.
-A.7 Additional Analysis of Attention
+A.8 Additional Analysis of Attention
 Figure 11: Learned per-instruction soft attention observed in the lower layers
@@
 In addition, in lower layers, we have observed attention weights to be softly distributed among tokens comprising each instruction (Figure 11), which suggests Nova initially models cross-relations among operation codes and operands in the first few layers, and later pools their summary representation into the [INST] token in the later layers.
+A.9 Limitations
+One limitation is that Nova is X86-specific, as we only collect X86 assembly corpus for pre-training.
+This design choice is mainly affected by two facts: (1) X86 assembly is used and explored in a wide range of binary tasks (Wang et al., 2022; Su et al., 2024; Xu et al., 2023a; Chen et al., 2022c) compared to other assembly languages, and (2) computation limitations.
+However, the proposed techniques are independent of X86 assembly.
+Low information density and compiler optimization are the common challenges of most assembly languages such as X86, ARM, and MIPS.
+The proposed techniques can be applied to the future development of multi-lingual assembly LLMs.
+Another potential limitation is the scale of models.
+We develop Nova-1.3B and Nova-6.7B and demonstrate their advancement in assembly code decompilation and encoding.
+Developing larger Nova models is promising, but remains as future work due to limited computing resources.

### v7

@@
 1 Introduction
-Binary code plays an irreplaceable role in the security domain, being the foundation of crucial tasks including vulnerability detection (Güler et al., 2019; Duan et al., 2020; Chen et al., 2022b), malware detection (Spensky et al., 2016; Aonzo et al., 2023; Xu et al., 2014), binary recovery (Su et al., 2024; Zhang et al., 2021; Chen et al., 2022c), and legacy software maintenance (Carbone et al., 2009; Carlini et al., 2015; Martin et al., 2010).
-For example, when performing tasks such as identifying attacks and malware, security analysts often only have access to assembly, i.e., the human-readable representation of binary code, which is extremely difficult to understand (Su et al., 2024; Zhang et al., 2021; Chen et al., 2022c).
-Thus, combined with the increasing sophistication of cybercrime that poses significant threats worldwide (e.g., cybercrime is predicted to cost the world \$10.5 trillion annually by 2025 (Sausalito, 2020)), effective binary analysis techniques are in high demand.
+Binary code plays an irreplaceable role in the security domain, being the foundation of crucial tasks including vulnerability detection (vul-motivate-1; vul-motivate-2; vul-motivate-3), malware detection (malware-motivate-1; malware-motivate-2; malware-motivate-3), binary recovery (codeart; osprey; dirty), and legacy software maintenance (legacy-motivate-1; legacy-motivate-2; legacy-motivate-3).
+For example, when performing tasks such as identifying attacks and malware, security analysts often only have access to assembly, i.e., the human-readable representation of binary code, which is extremely difficult to understand (codeart; osprey; dirty).
+Thus, combined with the increasing sophistication of cybercrime that poses significant threats worldwide (e.g., cybercrime is predicted to cost the world \$10.5 trillion annually by 2025 (cyber-attack)), effective binary analysis techniques are in high demand.
 Figure 1: Example that shows the semantics and diverse optimizations of assembly code.
-Large language models pre-trained on source code have brought improvement in various software development domains (Chen et al., 2022a; Liu et al., 2023a; Chen et al., 2023; Le et al., 2022; Jiang et al., 2023; Xia et al., 2023).
-However, these LLMs are not designed for or trained with assembly corpus, not achieving their full potential on binary code analysis tasks such as binary code similarity (Wang et al., 2022; Xu et al., 2023a), malware detection (Su et al., 2024), and binary code decompilation (Tan et al., 2024; Armengol-Estapé et al., 2024; Hosseini & Dolan-Gavitt, 2022).
-Existing work applying LLMs on assembly code mainly piggybacks on encoder-style LLMs (Wang et al., 2022; Su et al., 2024; Xu et al., 2023a), unable to benefit from the more extensive pre-training, updated architectures, scaling of state-of-the-art generative LLMs.
-Other work using generative LLMs for decompilation shows a low unit test passing rate of the decompiled programs (Tan et al., 2024; Armengol-Estapé et al., 2024).
+Large language models pre-trained on source code have brought improvement in various software development domains (related-generation-1; related-generation-2; related-generation-3; related-generation-4; related-apr-1; related-apr-2).
+However, these LLMs are not designed for or trained with assembly corpus, not achieving their full potential on binary code analysis tasks such as binary code similarity (jtrans; xu2023diemph), malware detection (codeart), and binary code decompilation (llm4decompile; slade; related-decompile-1).
+Existing work applying LLMs on assembly code mainly piggybacks on encoder-style LLMs (jtrans; codeart; xu2023diemph), unable to benefit from the more extensive pre-training, updated architectures, scaling of state-of-the-art generative LLMs.
+Other work using generative LLMs for decompilation shows a low unit test passing rate of the decompiled programs (llm4decompile; slade).
 The challenges of leveraging generative LLMs for assembly code are twofold.
@@
 Machine learning models are widely used in binary program analysis tasks.
-However, these models are typically designed for specific tasks such as binary code similarity detection (Pei et al., 2020; Xu et al., 2023a; Wang et al., 2022; Xu et al., 2017; Ding et al., 2019), variable name prediction (Chen et al., 2022c; Xu et al., 2023b; Zhang et al., 2021; He et al., 2018; Lacomis et al., 2020), binary code type inference (Pei et al., 2021), and so on (Chen et al., 2022d; Liu et al., 2023b; Hosseini & Dolan-Gavitt, 2022).
+However, these models are typically designed for specific tasks such as binary code similarity detection (pei2020trex; xu2023diemph; jtrans; related-bcsd-1; asm2vec), variable name prediction (dirty; xu2023lmpa; osprey; debin; dire), binary code type inference (pei2021stateformer), and so on (related-reverse-1; related-decompile-2; related-decompile-1).
 Recent techniques have started to pre-train foundation LLMs for binaries.
-CodeArt (Su et al., 2024) pre-trains encoder-style LLMs with a regularized attention design to better encode assembly code semantics.
-SLaDe (Armengol-Estapé et al., 2024) trains BART (Lewis et al., 2019) models on assembly.
-Meta LLMCompiler (Cummins et al., 2024) train CodeLlama models on LLVM IR to optimize binary code.
-LLM4Decompile (Tan et al., 2024) trains DeepSeekCoder with assembly for binary code decompilation.
+CodeArt (codeart) pre-trains encoder-style LLMs with a regularized attention design to better encode assembly code semantics.
+SLaDe (slade) trains BART (bart) models on assembly.
+Meta LLMCompiler (llm-compiler) train CodeLlama models on LLVM IR to optimize binary code.
+LLM4Decompile (llm4decompile) trains DeepSeekCoder with assembly for binary code decompilation.
 However, CodeArt does not generalize to generation tasks due to its encoder architecture.
@@
 2.2 Large Source-Code Models
-LLMs demonstrate promising results on many code-related tasks, such as code generation (Chen et al., 2022a; Liu et al., 2023a; Chen et al., 2023; Le et al., 2022; Yue et al., 2021; Chen et al., 2021; Nijkamp et al., 2022; Fried et al., 2023; Rozière et al., 2023; Guo et al., 2024; Lozhkov et al., 2024; Hui et al., 2024), bug fixing (Jiang et al., 2023; Xia et al., 2023) and vulnerability fixing (Wu et al., 2023; Steenhoek et al., 2023; He & Vechev, 2023).
-The advances in using LLMs are attributed to the knowledge learned from massive source code and natural language text in their training datasets (Touvron et al., 2023; OpenAI, 2023).
+LLMs demonstrate promising results on many code-related tasks, such as code generation (related-generation-1; related-generation-2; related-generation-3; related-generation-4; codet5; codex; codegen; incoder; codellama; deepseekcoder; starcoder2; qwencoder), bug fixing (related-apr-1; related-apr-2) and vulnerability fixing (related-vul-1; related-vul-2; related-vul-4).
+The advances in using LLMs are attributed to the knowledge learned from massive source code and natural language text in their training datasets (llama; openai2023gpt4).
 Nova is designed and trained for assembly, which has unique challenges such as low information density and diverse optimization.
 2.3 Attention Mechanism
-Standard self-attention is widely used in transformer architecture (Vaswani et al., 2017) to capture soft dependencies between tokens in the input.
-Many special attention mechanisms have been designed for better learning in various scenarios (Yang et al., 2016; Huang et al., 2024).
-LongCoder (Guo et al., 2023) combines window attention and global attention (attention sink (Xiao et al., 2024)) to handle long input of source code.
+Standard self-attention is widely used in transformer architecture (attention) to capture soft dependencies between tokens in the input.
+Many special attention mechanisms have been designed for better learning in various scenarios (hierarchical-attention; long-context-survey).
+LongCoder (longcoder) combines window attention and global attention (attention sink (streamingllm)) to handle long input of source code.
 We have shown that LongCoder’s window attention is less effective than Nova’s on assembly code.
-CAST (Shi et al., 2021) is a new neural architecture that splits the abstract syntax tree (AST) of source code into subtrees, encodes the subtrees, and aggregates to the final encoding.
-PA-former (Chai & Li, 2024) is a new neural architecture that constructs source code as pyramid input based on their AST structure and contains a pyramid attention mechanism to calculate the features in a hierarchical way.
-HierarchyNet (Nguyen et al., 2023) is a neural architecture that considers source code AST, data flow, and control flow graphs.
+CAST (cast) is a new neural architecture that splits the abstract syntax tree (AST) of source code into subtrees, encodes the subtrees, and aggregates to the final encoding.
+PA-former (pyramid-attention) is a new neural architecture that constructs source code as pyramid input based on their AST structure and contains a pyramid attention mechanism to calculate the features in a hierarchical way.
+HierarchyNet (hierarchy-net) is a neural architecture that considers source code AST, data flow, and control flow graphs.
 Similarly, it cannot be applied to assembly code.
@@
 Figure 2 presents the overall approach of Nova.
-We build Nova on top of foundation models for source code (Guo et al., 2024) to utilize their source code and natural language generation ability.
+We build Nova on top of foundation models for source code (deepseekcoder) to utilize their source code and natural language generation ability.
 We first collect large assembly corpora (Section 3.1).
@@
 In this paper, we focus on X86-64 assembly functions for C programs. Yet, Nova’s approach is generalizable to other assembly languages such as ARM assembly.
-We derive our X86-64 assembly functions dataset from two source code corpora: C functions in The-Stack (Li et al., 2023) and AnghaBench (da Silva et al., 2021).
+We derive our X86-64 assembly functions dataset from two source code corpora: C functions in The-Stack (starcoder) and AnghaBench (anghabench).
 We compile the C programs into executables using gcc with different optimization levels (i.e., O0, O1, O2 and O3), strip the executables to remove debug information, and disassemble them into X86-64 assembly code using objdump.
@@
 Specifically, Nova learns the assembly code in an hierarchical way by providing a modified attention mask.
-Different from standard token-level attentions (Vaswani et al., 2017; Radford & Narasimhan, 2018; Radford et al., 2019; Brown et al., 2020), our hierarchical self-attention contains three categories at different levels of granularity.
+Different from standard token-level attentions (attention; gpt-1; gpt-2; gpt-3), our hierarchical self-attention contains three categories at different levels of granularity.
 Figure 3: Design of Nova’s hierarchical attention for assembly code
@@
 The syntax gap between assembly code and source code, and syntax diversity between differently-optimized assembly code make LLMs struggle to distinguish the semantics behind the syntax.
-Nova adopts contrastive learning technique (Gao et al., 2021) during pre-training to train LLMs to encode assembly code meaningfully w.r.t semantics.
-The standard pre-training objective is language modeling by minimizing the negative likelihood of code in the pre-training corpus (Radford & Narasimhan, 2018), notated as $L_{lm}$.
+Nova adopts contrastive learning technique (simcse) during pre-training to train LLMs to encode assembly code meaningfully w.r.t semantics.
+The standard pre-training objective is language modeling by minimizing the negative likelihood of code in the pre-training corpus (gpt-1), notated as $L_{lm}$.
 In addition, Nova is pre-trained with two new objectives, $L_{fcl}$ for functionality contrastive learning and $L_{ocl}$ for optimization contrastive learning.
@@
 For assembly, we use the average of all the “[INST]” tokens’ embedding as the embedding of the assembly function, as each “[INST]” token is supposed to capture the semantics of that instruction by the design of our hierarchical self-attention.
-Let $e^{s}_{f}$ be the embedding of function $f$ in $s$ form ($s=-1$ for source code, and $s\in[0,1,2,3]$ for O0 to O3 optimized assembly).
-For simplicity, let $S=[-1,0,1,2,3]$ be the domain of $s$.
+Let $e^{s}_{f}$ be the embedding of function $f$ in $s$ form (s=1s=-1 for source code, and $s\in[0,1,2,3]$ for O0 to O3 optimized assembly).
+For simplicity, let S=[1,0,1,2,3]S=[-1,0,1,2,3] be the domain of $s$.
 Functionality CL optimizes Nova’s embeddings to satisfy the constraint:
-[…]
-[…]
+$\forall f_{i}\in F,\;\mathop{max}\limits_{s,t\in S}(\mathop{d}(e^{s}_{f_{i}},e^{t}_{f_{i}}))<\mathop{min}\limits_{\begin{subarray}{c}s,t\in S\\ f_{j}\neq f_{i}\in F\end{subarray}}(\mathop{d}(e^{s}_{f_{i}},e^{t}_{f_{j}}))$
+, where $\mathop{d}$ calculates the $l_{2}$ distance between two embeddings and $F$ is the full set of functions in the training corpus.
 The embeddings of a batch of functions, each represented in two different forms, can be optimized to satisfy these constraints.
@@
 3.4 Task 1: Binary Code Decompilation
-[…]
+Binary code decompilation (BCD) helps developers to understand binary code by recovering binary code into more readable high-level source code (e.g., C programs) (coda; neutron; slade; llm4decompile).
+The input to the model for BCD is formatted as an instruction prompt (notated by $\mathbf{p}$): # This is the assembly code with {opt} optimization: $\mathbf{asm}$}, where “opt” is the optimization-level applied to the assembly and $\mathbf{asm}$” is the assembly code to decompile.
+Nova is fine-tuned to generate the expected source code function $\mathbf{src}$ following the instruction prompt.
+The fine-tuning objective is minimizing the loss: $L_{bcd}=-\mathop{log}\mathop{P}(\mathbf{src}|\mathbf{p})$.
 3.5 Task 2: Binary Code Similarity Detection
-Binary code similarity detection (BCSD) aims to measure the similarity between two binary code snippets (Wang et al., 2022; Su et al., 2024), which is the foundation of various applications such as plagiarism detection (Luo et al., 2014; Sæbjørnsen et al., 2009) and vulnerability detection (David & Yahav, 2014; David et al., 2018; 2017; 2016).
+Binary code similarity detection (BCSD) aims to measure the similarity between two binary code snippets (jtrans; codeart), which is the foundation of various applications such as plagiarism detection (plagiarism-1; plagiarism-2) and vulnerability detection (binary-vul-1; binary-vul-2; binary-vul-3; binary-vul-4).
 A widely used setting is taking a query assembly of the function $f^{q}$ that is compiled with one optimization level (denoted by $s$), and a pool of candidate assembly of $K$ (e.g., 50, 100, etc.) functions (notated by $f^{p}_{i},$ $1\leq i\leq K$) compiled with a different optimization level (denoted by $t\neq s$).
@@
 The learning objective is as follows:
-[…]
-We follow previous work (Su et al., 2024) to let $s$ be O0-assembly and $t$ be O3-assembly, which is the hardest setting.
+$L_{BCSD}=-\mathop{log}\sum\limits_{1\leq i\leq K}^{f^{q}:=f^{p}_{i}}\left(1-\frac{\mathop{exp}\left(\mathop{d}(e^{s}_{f^{q}},e^{t}_{f^{p}_{i}})\right)}{\sum_{1\leq j\leq K}\mathop{exp}\Big(\mathop{d}(e^{s}_{f^{q}},e^{t}_{f^{p}_{j}})\Big)}\right)$
+We follow previous work (codeart) to let $s$ be O0-assembly and $t$ be O3-assembly, which is the hardest setting.
 4 Experimental Setup
@@
 We use the C and X86-64 assembly functions collected from AnghaBench and The-Stack for pre-training.
-We pre-train Nova starting from DeepSeek-Coder (Guo et al., 2024), and the hierarchical attention is applied on half of the attention heads to balance between its effectiveness and the existing knowledge in the standard attention layers (Justified in Appendix A.3).
+We pre-train Nova starting from DeepSeek-Coder (deepseekcoder), and the hierarchical attention is applied on half of the attention heads to balance between its effectiveness and the existing knowledge in the standard attention layers (Justified in Appendix A.3).
 Nova is pre-trained with language modeling for one epoch, followed by contrastive learning objectives for another epoch.
@@
 Training Data: We sample (due to computation resource limitation) 2.16M assembly-to-source-code pairs (0.338B tokens) from the pre-training corpus to build the BCD fine-tuning data.
-Test Data: We use HumanEval-Decompile (Tan et al., 2024) as the test benchmark, which was not used in training.
-HumanEval-Decompile is derived from the C language adaptation of the HumanEval (Chen et al., 2021) benchmark and contains 164 C functions, each compiled with O0O3 optimization flags and disassembled into X86-64 assembly.
-Baselines: Nova is compared with existing SOTA LLM4Decompile (Tan et al., 2024), Meta LLMCompiler (Cummins et al., 2024), other open-sourced general code LLMs (Rozière et al., 2023; Lozhkov et al., 2024; Guo et al., 2024; Hui et al., 2024; Mishra et al., 2024; Guo et al., 2023), and commercial LLMs GPT-3.5-Turbo and GPT-4o).
+Test Data: We use HumanEval-Decompile (llm4decompile) as the test benchmark, which was not used in training.
+HumanEval-Decompile is derived from the C language adaptation of the HumanEval (codex) benchmark and contains 164 C functions, each compiled with O0O3 optimization flags and disassembled into X86-64 assembly.
+Baselines: Nova is compared with existing SOTA LLM4Decompile (llm4decompile), Meta LLMCompiler (llm-compiler), other open-sourced general code LLMs (codellama; starcoder2; deepseekcoder; qwencoder; granite; longcoder), and commercial LLMs GPT-3.5-Turbo and GPT-4o).
 LLM4Decompile trains DeepSeekCoder using the same AnghaBench corpus for binary decompilation.
 Meta LLMCompiler trains CodeLlama models using LLVM IRs, X86, and ARM assembly code to optimize and translate binary code.
-Evaluation: We let each model sample 20 decompilations per assembly function, using the temperature of 0.2 and top_p of 0.95 (Chen et al., 2021).
-Except for LLM4Decompile and Nova that are fine-tuned for binary code decompilation, we provide all other baselines with three-shot examples for few-shot learning (Brown et al., 2020).
+Evaluation: We let each model sample 20 decompilations per assembly function, using the temperature of 0.2 and top_p of 0.95 (codex).
+Except for LLM4Decompile and Nova that are fine-tuned for binary code decompilation, we provide all other baselines with three-shot examples for few-shot learning (gpt-3).
 The decompilations are executed with the test cases to verify the functional correctness.
-Finally, Pass@1 and Pass@10 (Chen et al., 2021) are reported.
+Finally, Pass@1 and Pass@10 (codex) are reported.
 4.3 Fine-Tuning for Binary Code Similarity Detection
-Training Data: To compare Nova with existing works on BCSD fairly (Wang et al., 2022; Su et al., 2024), we use the same dataset, BinaryCorp-3M (Wang et al., 2022), as the fine-tuning data for BCSD, which contains the O0 and O3 assembly of 224,606 functions.
-Test Data: Following existing work (Su et al., 2024; Xu et al., 2023a), we use real-world benchmarks, Binutils, Curl, ImageMagick, SQLite, OpenSSL, and Putty, as the test benchmarks, which are nonexistent in the training data.
-Baselines: Nova is compared with jTrans (Wang et al., 2022), DiEmph (Xu et al., 2023a) and CodeArt (Su et al., 2024).
-jTrans is a Transformer (Vaswani et al., 2017) encoder trained on binaries with masked token prediction and jump target prediction tasks.
+Training Data: To compare Nova with existing works on BCSD fairly (jtrans; codeart), we use the same dataset, BinaryCorp-3M (jtrans), as the fine-tuning data for BCSD, which contains the O0 and O3 assembly of 224,606 functions.
+Test Data: Following existing work (codeart; xu2023diemph), we use real-world benchmarks, Binutils, Curl, ImageMagick, SQLite, OpenSSL, and Putty, as the test benchmarks, which are nonexistent in the training data.
+Baselines: Nova is compared with jTrans (jtrans), DiEmph (xu2023diemph) and CodeArt (codeart).
+jTrans is a Transformer (attention) encoder trained on binaries with masked token prediction and jump target prediction tasks.
 DiEmph uses an instruction deemphasis technique to prevent the model from learning instruction distribution biases introduced by compilers.
@@
 Morgan AI Faculty Research Awards.
+This work also used Anvil at Purdue University through allocation CIS240304 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program (anvil), which is supported by NSF grants 2138259, 2138286, 2138307, 2137603, and 2138296.
 Any opinions, findings, and conclusions in this paper are those of the authors only and do not necessarily reflect the views of our sponsors.
@@
 During the training using the contrastive learning objectives, and the fine-tuning of BCSD, we use the average of [INST] tokens’ last layer hidden states to represent the embedding of a binary function.
-For the evaluation of BCSD, we reuse the framework provided by CodeArt Su et al.
-(2024) to evaluate the binary code similarity detection results once Nova produces the embeddings for functions in the test dataset.
+For the evaluation of BCSD, we reuse the framework provided by CodeArt codeart to evaluate the binary code similarity detection results once Nova produces the embeddings for functions in the test dataset.
 For each one of the $K$ functions, the O0 assembly function is used as the query to calculate the cosine similarity between its embedding and the embeddings of all the $K$ O3 assembly functions.
@@
 A.6 Binary Code Similarity Detection Ablation Study
-Table $15$ show the detailed ablation study results of BCSD. Nova wins on the most benchmarks when $K=100$ or $500$, and ties with Nova-HA when $K=50$, or $200$.
+Table 15151515 show the detailed ablation study results of BCSD. Nova wins on the most benchmarks when $K=100$ or $500$, and ties with Nova-HA when $K=50$, or $200$.
 Figure 9: PCA of embeddings calculated by Nova-CL-HA, Nova-HA, and Nova.
@@
 opcodes and operands, shown in Figure 10 (d), attentions are paid to “push”, “mov”, etc.).
-Quantitatively, we have determined broad attention accounts for as much as 30% of all attention in standard heads, especially in layers 1-8 (consistent with the findings of (Clark et al., 2019)), whereas in Nova’s hierarchical attention, no more than 5% or all attention is allocated to each instruction segment.
+Quantitatively, we have determined broad attention accounts for as much as 30% of all attention in standard heads, especially in layers 1-8 (consistent with the findings of (explain-bert)), whereas in Nova’s hierarchical attention, no more than 5% or all attention is allocated to each instruction segment.
 This validates our goal of learning instruction-aware hierarchical attention in Nova.
@@
 One limitation is that Nova is X86-specific, as we only collect X86 assembly corpus for pre-training.
-This design choice is mainly affected by two facts: (1) X86 assembly is used and explored in a wide range of binary tasks (Wang et al., 2022; Su et al., 2024; Xu et al., 2023a; Chen et al., 2022c) compared to other assembly languages, and (2) computation limitations.
+This design choice is mainly affected by two facts: (1) X86 assembly is used and explored in a wide range of binary tasks (jtrans; codeart; xu2023diemph; dirty) compared to other assembly languages, and (2) computation limitations.
 However, the proposed techniques are independent of X86 assembly.